### 17 Mar 2026 v2

# Robust Optimal Strategies for Early Liquidation in Financial Systems

Abstract
We study the problem of asset liquidation in financial systems. During financial crises, asset liquidation is often inevitable but can lead to substantial losses if a significant amount of illiquid assets are sold simultaneously at depressed prices—a phenomenon known as price impact. To tackle this challenge, we consider a two-period liquidation model that allows for early liquidation prior to clearing, thereby mitigating price impact at clearing, and we develop a worst-case approach to solve the decision-making problem on the optimal size of early liquidation. Specifically, we propose a robust optimal strategy—a tractable liquidation approach that maximizes the worst-case value of liquid assets at clearing, taking into account the uncertainty of other banks’ early liquidation decisions. We derive a (semi-)closed-form representation of this strategy in a practical scenario involving permanent price impact and analyze its sensitivity to that impact’s magnitude. We further identify its closed-form expression in another practical scenario featuring interbank exposures. Our findings, although built upon a stylized model, offer valuable guidelines for developing robust liquidation strategies that mitigate losses resulting from asset liquidation.
Keywords: Systemic risk, Fire sales, Strategic liquidation, Worst-case approach, Permanent price impact, Interbank exposures
1.  Introduction
The global financial crisis in 2007-2008 provoked intense discussions on the factors that threatened the stability of the financial system. In the majority of relevant studies, two channels of risk contagion have been widely acknowledged as significant factors contributing to this instability. One such channel involves direct debt exposures among financial institutions. When an institution fails to meet debt obligations due to adverse economic shocks, its creditor institutions experience financial distress due to the resulting repayment shortfall (Eisenberg and Noe, 2001). Furthermore, during the crisis of asset erosion, financial institutions with inadequate liquid assets are forced to sell off substantial illiquid assets to satisfy their debt claims. Such liquidation overwhelms shallow market demand and triggers sharp declines in asset prices. This phenomenon, termed price impact, imposes significant losses on institutions holding the same illiquid assets, even without direct debt exposures among them. These fire sales constitute another channel of risk transmission and are recognized as a salient ingredient of the crisis (Clerc et al., 2016). Although the first channel has been studied more extensively in the literature, the second channel has gained increasing attention in recent years, particularly as financial institutions have actively reduced their debt exposures to other banks since the said crisis.
Our study departs from a conventional assumption in the existing literature that liquidation is conducted only at clearing. This assumption stems from the observation that liquidity-constrained institutions tend to favor borrowing over asset sales to avoid the prohibitive costs of liquidation. However, since borrowing capacity is not always guaranteed, relying solely on this option exposes banks to future funding shortages, particularly during systemic crises when credit conditions deteriorate (Acharya et al., 2011). In this case, preemptive liquidation prior to clearing becomes a rational solution. Then, this naturally raises a critical question: what volume of illiquid assets should an institution sell early to ensure its solvency? The answer is not immediately evident as liquidation outcomes are driven by the interplay of institutional liquidation decisions that jointly determine sale prices and are constrained by repayment requirements.
This study is the first to propose a solution to address the aforementioned issue, particularly focusing on two distinct practical settings: permanent price impact induced by early liquidation and direct loss contagion through interbank exposures. Specifically, our main contributions are summarized as follows:
  1. 1.
    Design of a novel decision-theoretic framework for early liquidation. We develop a two-period liquidation model that offers flexibility in allocating liquidation volume over time and endogenizes the interdependence of institutional liquidation decisions. Based on this model, we adopt a worst-case approach to optimize an individual bank’s early liquidation strategy by formulating a maximin problem that maximizes its liquidation proceeds under the worst-case risk contagion arising from counterparties’ strategies. This formulation serves as the analytical foundation for quantitatively evaluating early liquidation decisions.
  2. 2.
    Tractable relaxation for the maximin liquidation problem. A critical challenge of the said maximin problem is its analytical intractability, which stems from a multi-layer optimization structure driven by implicit strategic interactions among banks. To overcome this, we propose a relaxed version of the problem, which replaces the liquidation size at clearing by its worst-case upper bound, and establish its validity as a robust surrogate for the original formulation.
  3. 3.
    Analytical solution to the relaxed problem. We demonstrate that the relaxed problem yields a unique solution—hereafter referred to as robust optimal strategy—that can be expressed in (semi-)closed form, even in complex scenarios involving permanent price impacts or interbank exposures. This analytical form of the strategy not only reduces the computational burden relative to the original problem by circumventing the need for exhaustive search over high-dimensional spaces but also facilitates effective decision support by providing practitioners with a robust tool to secure liquidity positions during early liquidation, even without observable data on counterparty actions.
  4. 4.
    Practical insights into strategic behaviors. Our analysis demystifies the strategic drivers of fire sales. We show that while early liquidation is theoretically capped at half the total liquidation volume required in the absence of early action, the incentive to sell early depends on a trade-off between immediate competition and future price risks. In particular, we identify a critical tipping point governed by market conditions. Initially, a higher permanent price impact (i.e., slower price recovery) prompts counterparties to sell early in the worst-case scenario, leading an individual bank is to reduce its own early liquidation to avoid crowded trades. Beyond this threshold, however, the magnified risk of a future price collapse dominates, paradoxically driving an increase in the robust optimal early liquidation size.
The remainder of the paper is organized as follows. Section 2 provides a review of the related literature. In Section 3, we identify key issues and motivate our research questions through an illustrative example. Section 4 derives the robust optimal strategy for the case with permanent price impact and analyzes its sensitivity to the magnitude of the impact. In Section 5, we perform a similar analysis for the case with interbank liabilities. Section 6 concludes the paper. All proofs of the theoretical results can be found in the appendix.
2.  Literature Review
Impact of fire sales on systemic risk. Our research builds upon the literature on risk contagion via fire sales of illiquid assets, initiated by Cifuentes et al. (2005). They investigate the equilibrium of liquidation prices in financial systems where banks are forced to liquidate assets to satisfy regulatory constraints. Chen et al. (2016) formulate this equilibrium as an optimization problem and introduce a “liquidity amplifier” via sensitivity analysis, which quantifies how illiquidity intensifies fire-sale effects, thereby offering policy implications for intervention. In addition, Amini et al. (2016) establish sufficient conditions for the uniqueness of the equilibrium liquidation price, while Weber and Weske (2017) characterize the liquidation price in more complex settings where bankruptcy costs and cross-holdings are taken into account. These papers mainly focus on the assessment of systemic risk rather than the decision-making in liquidation and assume that liquidation takes place only at clearing. By contrast, we accommodate liquidation that occurs before clearing and offer novel insights into strategic liquidation decisions. A distinct line of work, pioneered by Greenwood et al. (2015), leverages correlations in banks’ asset exposures to measure fire-sale spillover effects. Subsequent studies have enriched this paradigm by modeling the dynamic process of fire sales (Duarte and Eisenbach, 2021) and incorporating partial information about asset holdings (Pang and Veraart, 2023). Although their approaches open up new directions, our modeling focus remains on extending the framework of Cifuentes et al. (2005), leaving the other models for further research.
Strategic liquidation in financial systems. Our work contributes to the growing literature of strategic liquidation under systemic risk. Feinstein (2017) studies a financial system with multiple illiquid assets in which firms strategically choose asset sales to maximize their own valuations, and establishes the existence of equilibrium liquidation strategies. To analyze risk contagion arising from strategic liquidation, Braouezec and Wagalath (2019) consider a static game in which banks maximize their expected profits subject to regulatory capital requirements and characterize the resulting Nash equilibrium. Bichuch and Feinstein (2019) formulate an optimization problem in which borrowing and liquidation decisions are jointly determined to minimize total financing costs. While these studies analyze strategic liquidation through ex-post equilibrium outcomes of institutional interactions assuming rational behaviors of institutions, our work focuses on the ex-ante liquidation decision of an individual institution adopting a worst-case approach to the characterization of early liquidation strategies. [^1] It is essential to distinguish the notion of strategic liquidation considered here from the classical literature on optimal execution (e.g., Bertsimas and Lo, 1998; Gatheral and Schied, 2013), which focuses on minimizing trading costs when unwinding a fixed asset position over a given period under exogenous market illiquidity. In our systemic context, by contrast, illiquidity is not merely a frictional cost but an endogenous channel of risk contagion that amplifies price declines and threatens systemic stability.
The role of permanent price impact in financial systems. Our research is closely related to studies on liquidation-induced systemic risk in the presence of permanent price impact, where price declines caused by liquidation do not immediately reverse and losses from liquidation accumulate over time. Cont and Schaanning (2019) incorporate such cumulative impacts to empirically measure indirect contagion from fire sales empirically. Capponi et al. (2020) examine a dynamic setting where persistent price impacts can trigger a self-reinforcing cycle of mutual fund redemptions and further liquidations. In the context of collateralized systems, Ghamami et al. (2022) and Pang and Veraart (2025) model clearing processes in which collateral is liquidated in multiple rounds, with price impacts from early sales carrying over entirely to subsequent rounds. While these works assume fully permanent price impact, implying that depressed prices do not recover, our work allows for partial price recovery. Furthermore, we examine robust liquidation decisions under permanent price impact from an individual decision-making perspective, whereas the aforementioned studies primarily focus on measuring systemic risk or elucidating contagion mechanisms.
Contagion through interbank exposures. A large body of literature on systemic risk investigates the impact of interbank obligations. The seminal framework of Eisenberg and Noe (2001) models interbank repayments using a fixed-point characterization. Subsequent research has extended this framework along several dimensions while retaining its core analytical structure. For instance, Rogers and Veraart (2013) incorporate bankruptcy costs into the framework and explore incentives for consortium-based bailouts. Acemoglu et al. (2015) and Capponi et al. (2016) investigate how network topology and concentration affect shock transmission, respectively. Kusnetsov and Veraart (2019) and Veraart (2020) extend the framework to accommodate multiple maturities and quantify distress versus default contagion, respectively. Furthermore, this approach has been applied to diverse areas such as systemic stress scenario selection (Ahn et al., 2023), debt valuation (Barucca et al., 2020; Banerjee and Feinstein, 2022), intervention policies (Ahn and Kim, 2019; Bernard et al., 2022), and multilateral netting effects (Ahn, 2020; Amini et al., 2020). We also incorporate interbank exposures as an additional risk contagion channel. However, in contrast to the aforementioned works that analyze systemic risk from a regulatory perspective, our work examines the impact of this contagion channel on early liquidation decisions from the standpoint of an individual bank.
3.  Research Questions
We motivate our research questions via an illustrative example—a special case of our main problems.
An illustrative example. Consider a system of $m$ banks operating over two periods ($t=1,2$), where liabilities are cleared at the end of the time horizon. Each bank $i$ possesses liquid assets valued at $e_{i}$ and holds $z_{i}$ units of illiquid assets. If the liquid assets prove insufficient to meet liabilities $L_{i}$ (i.e., $e_{i}<L_{i}$), the bank must cover the shortfall by selling off its illiquid holdings at a depressed price—a process known as fire sales. In contrast to the standard literature that assumes no early fire sales, we allow each bank $i$ to strategically liquidate its illiquid holdings $x_{i}$ at $t=1$, prior to clearing at $t=2$, in order to mitigate price impacts at clearing and sell a smaller amount of illiquid assets than in the scenario without early liquidation. Given a realization ${\bf x}=(x_{1},\ldots,x_{m})^{\top}$ of all banks’ early liquidation decisions, let $\psi_{i}({\bf x})$ denote the amount of illiquid assets bank $i$ must sell at clearing to fulfill its debt obligation for each $i=1,\ldots,m$. Then, by defining ${\bf x}_{-i}\coloneqq(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{m})^{\top}$, bank $i$’s liquid asset value at clearing becomes:
$\phi_{i}(x_{i},{\bf x}_{-i})\coloneqq e_{i}+x_{i}Q\left(\sum_{j=1}^{m}x_{j}\right)+\psi_{i}({\bf x})Q\left(\sum_{j=1}^{m}\psi_{j}({\bf x})\right),$ (1)
where $Q(\cdot)$ maps the total supply of illiquid assets to their price, and thus, the second and third terms represent proceeds from asset liquidation at $t=1$ and $t=2$, respectively.
Robust decision-making. In this example, each bank $i$ would seek to determine its strategy $x_{i}$ that optimizes $\phi_{i}(x_{i},{\bf x}_{-i})$ while maintaining robustness against other banks’ strategies ${\bf x}_{-i}$, as these are practically unobservable. However, as well documented in the literature (Amini et al., 2016) and as we shall demonstrate later, the mapping $\psi_{j}(\cdot)$ is characterized by a fixed-point equation involving nonconvex functions and is hence analytically intractable. A natural approach to this issue would be to replace $\psi_{j}({\bf x})$ by its surrogate $\psi_{j}({\bf 0})-x_{j}$ for all $j$, where $\bf 0$} denotes the zero vector. [^2] The early liquidation size $x_{j}$ is capped at $\psi_{j}({\bf 0})$ for all $j=1,\ldots,m$. This is because, as noted earlier, the goal of early liquidation is to reduce the amount of illiquid assets sold and the losses from liquidation compared to the case without it. Similarly, $\psi_{j}({\bf x})$ is capped at $\psi_{j}({\bf 0})-x_{j}$. This accounts for the worst-case scenario in which banks eventually sell off the same amount of illiquid assets as they would in the absence of early liquidation. Then, the said decision-making problem of bank $i$ can be formulated as follows:
$\max_{x_{i}}\min_{{\bf x}_{-i}}$ $e_{i}+x_{i}Q\left(\sum_{j=1}^{m}x_{j}\right)+\left(\bar{x}_{i}-x_{i}\right)Q\left(\sum_{j=1}^{m}(\bar{x}_{j}-x_{j})\right)$ (2)
s.t. $0\leq x_{j}\leq\bar{x}_{j}~~\text{for}~j=0,1,\ldots,m,$
where $\bar{x}_{j}\coloneqq\psi_{j}({\bf 0})$ for all $j=1,\ldots,m$. We refer to the optimal solution to the outer problem of (2), denoted by $x^{*}_{i}$, as bank $i$’s robust optimal strategy.
Solving (2) is surprisingly simple. By letting $s_{i}=\sum_{j\neq i}x_{j}$ and $\bar{s}_{i}=\sum_{j\neq i}\bar{x}_{j}$, it can be recast as $\max_{x_{i}\in[0,\bar{x}_{i}]}\min_{s_{i}\in[0,\bar{s}_{i}]}g_{i}(x_{i},s_{i})$, where $g_{i}(x,s)\coloneqq e_{i}+xQ(x+s)+(\bar{x}_{i}-x)Q(\bar{x}_{i}+\bar{s}_{i}-(x+s_{i}))$. Under the common assumptions on $Q(\cdot)$ in the literature (see Assumption 1), $g_{i}(x,s)$ is concave in $x$ and convex in $s$. Therefore, the minimax theorem (Rockafellar, 2015, Theorem 36.3) holds: $\max_{x_{i}\in[0,\bar{x}_{i}]}\min_{s_{i}\in[0,\bar{s}_{i}]}g_{i}(x_{i},s_{i})=\min_{s_{i}\in[0,\bar{s}_{i}]}\max_{x_{i}\in[0,\bar{x}_{i}]}g_{i}(x_{i},s_{i})$. Furthermore, $h_{i}$ satisfies the following point symmetry: $g_{i}(x_{i},s_{i})=g_{i}(\bar{x}_{i}-x_{i},\bar{s}_{i}-s_{i})$ for any $x_{i}\in[0,\bar{x}_{i}]$} and $s_{i}\in[0,\bar{s}_{i}]$}. Accordingly, $(\bar{x}_{i}/2,\bar{s}_{i}/2)$ is the saddle point of $h_{i}$, implying that
$x^{*}_{i}=\frac{\bar{x}_{i}}{2}.$ (3)
Main issues and questions. The simplicity of identifying the robust optimal strategy, however, is limited to the special case presented above. When practical considerations are incorporated, finding the strategy becomes significantly more complex. Specifically, this paper focuses on two distinct extensions prevalent in the literature: incorporating permanent price impact and including interbank exposures. In the former case, early liquidation exerts downward pressure on the asset price at clearing, which causes the associated version of $h_{i}$ to no longer exhibit point symmetry. In the latter case, interbank repayments must be factored into the value of liquid assets at clearing. This makes the corresponding version of $h_{i}$ dependent on the intricate interplay of liquidation strategies and payment requirements, thereby destroying its concave-convexity and precluding the application of the minimax theorem. Consequently, our primary research question is:
  • Q1.
    Can we characterize (semi-)closed-form expressions of robust optimal strategies in these two cases?
Additionally, we investigate the following question:
  • Q2.
    Is the idea of replacing $\psi_{j}({\bf x})$ with a surrogate $\bar{x}_{j}-x_{j}$ valid for identifying a good liquidation strategy?
We examine these questions in the subsequent sections.
Further notation. All vectors are column vectors and denoted by bold symbols, e.g., ${\bf u}=(u_{1},\dots,u_{d})^{\top}\in\mathbb{R}^{d}$. We use $[m]$ and $\bf 1$} to represent the set $\{1,\dots,m\}$ and the vector of ones in a suitable dimension, respectively. For any two vectors ${\bf u},{\bf v}\in\mathbb{R}^{d}$, ${\bf u}\leq{\bf v}$ means an entry-wise inequality, ${\bf u},{\bf v}]=[u_{1},v_{1}]\times\cdots\times[u_{d},v_{d}]$, and ${\bf u}\wedge{\bf v}=(\min\{u_{1},v_{1}\},\dots,\min\{u_{d},v_{d}\})^{\top}$. For any matrix $\bf M$}, we denote its $i$-th column by ${\bf M}_{i}$. Finally, for any $x\in\mathbb{R}$, $x^{+}\coloneqq\max\{x,0\}$.
4.  Robust Optimal Strategy with Permanent Price Impact
It is widely known in the literature that price impacts are often persistent; buyers observe the depressed post-liquidation price and anticipate a slow recovery, preventing immediate reversion. In this section, we analyze the impact of this permanent price impact on the robust optimal strategy by characterizing it under different degrees of permanent price impact.
4.1.  Model Setup
We detail our full modeling framework, elaborating on technical aspects omitted from the illustrative example and demonstrating how permanent price impact is incorporated into that example.
Balance sheet information. We consider a financial system comprised of $m$ banks, indexed by $1,\dots,m$, that operates over two periods ($t=1,2$). Each bank $i$’s balance sheet consists of the following four components:
  • $e_{i}>0$ stands for the “value” of its liquid assets;
  • $z_{i}\geq 0$ represents the “volume” of its illiquid assets, with a nominal unit price of PP; [^3] We adopt the standard assumption in the literature that all banks hold the same type of illiquid assets, which serves as an approximation of multiple correlated assets (Weber and Weske, 2017; Braouezec and Wagalath, 2019).
  • $L_{i}>0$ denotes its total liabilities to creditors;
  • $w_{i}\coloneqq e_{i}+z_{i}P-L_{i}$ means its initial net worth before liquidation.
We assume that these components remain constant throughout the two periods and that liquid assets (e.g., cash, central bank reserves, and short-term government securities) can be sold and transferred at their nominal value without any cost. To exclude trivial cases where banks are not required to liquidate their illiquid assets, it is assumed that $e_{i}<L_{i}$ for all $i\in[m]$. The associated balance sheet is described in Table 1.
Table 1: Initial balance sheet of bank $i$ before liquidation
Assets Liabilities and Net Worth
Liquid assets $e_{i}$ Liabilities $L_{i}$
Illiquid assets $z_{i}$ with price $P$ Net worth $w_{i}$
Illiquid asset price modeling. Illiquid assets (e.g., loans, mortgages, and real estate holdings) are, however, difficult to sell without a significant loss of value. In the related literature (Cifuentes et al., 2005; Chen et al., 2016), this loss is typically modeled by the so-called inverse demand function $Q:[0,\sum_{j=1}^{m}z_{j}]\rightarrow\mathbb{R}_{+}$, a decreasing function that maps the aggregate supply of illiquid assets across all banks to their price. Throughout this paper, we impose several regularity conditions on the function $Q$, described below:
Assumption 1.
The inverse demand function $Q$ is twice continuously differentiable and satisfies the following five conditions: (i) $Q(0)=P$; (ii) $Q^{\prime}(x)<0$; (iii) $(xQ(x))^{\prime}>0$; (iv) $Q^{\prime\prime}(x)\geq 0$; and (v) $(xQ^{\prime}(x))^{\prime}<0$.
The first condition states that illiquid assets are valued at their nominal price in the absence of sales. Conditions (ii) and (iii) reflect the observation that as the supply of illiquid assets increases, their price decreases while the total value of these assets in the market grows. The fourth condition imposes a diminishing marginal effect on the price of illiquid assets with respect to their size. Condition (v) ensures that the total value of illiquid assets in the market is concave in their size. Note that the above conditions (or their variants) are commonly used in the literature (Amini et al., 2016; Bichuch and Feinstein, 2019). Examples of inverse demand functions satisfying Assumption 1 include Q(x)=PaxQ(x)=P-ax with $0<a<P/(2\sum_{k=1}^{m}z_{k})$, $Q(x)=P\exp(-ax)$ with $0<a<1/\sum_{k=1}^{m}z_{k}$, and $Q(x)=P(\beta x+1)^{-a}$ with $0<a<1/(\beta\sum_{k=1}^{m}z_{k})$ and $\beta>0$.
Early liquidation and permanent price impact. We assume that all banks clear their debts at the end of the time horizon (i.e., $t=2$), consistent with prior studies (e.g., Cifuentes et al., 2005; Chen et al., 2016; Amini et al., 2016; Ghamami et al., 2022). As alluded to earlier, a salient feature of our modeling framework lies in the option for strategic early liquidation at $t=1$ to mitigate liquidation losses and to sell a smaller volume of illiquid assets, which is in stark contrast to most existing models that allow liquidation only at clearing. Suppose that each bank $i$ chooses to sell off $x_{i}$ units of illiquid assets at $t=1$. Then, $Q(\sum_{k=1}^{m}x_{k})$ becomes the market price of illiquid assets at $t=1$ and is strictly smaller than the nominal price $P$ by Assumption 1.
To account for permanent price impact, we consider a situation where the price impact generated at $t=1$ persists into $t=2$. Specifically, the asset price at clearing is governed by another inverse demand function of the asset supply at $t=2$, defined as $u\mapsto Q(\gamma\sum_{k=1}^{m}x_{k}+u)$, where the parameter $\gamma\in[0,1]$ represents the proportion of the early liquidation volume that continues to depress prices at $t=2$. That is, a larger $\gamma$ implies slower price recovery. The limiting case $\gamma=0$ corresponds to fully transient impact, where the initial price at $t=2$ recovers to the nominal price $P$, as in the example in Section 3. Conversely, $\gamma=1$ signifies fully permanent price impact, where the opening price at $t=2$ remains at the post-liquidation price at $t=1$; this latter case coincides with the model in Ghamami et al. (2022). While existing research commonly focuses on either of these two extreme cases, our approach generalizes the framework of Ghamami et al. (2022) by allowing for “partially” permanent price impact via the parameter $\gamma$.
Clearing process at $t=2$. Following the early liquidation at $t=1$, each bank $i\in[m]$ retains $z_{i}-x_{i}$ units of illiquid assets, while its liquid asset value becomes $e_{i}+x_{i}Q(\sum_{k=1}^{m}x_{k})$. At the clearing stage, given a price $p$, each bank $i$ attempts to cover its liabilities by selling $p^{-1}(L_{i}-e_{i}-x_{i}Q(\sum_{k=1}^{m}x_{k}))^{+}$ units of illiquid assets. If the bank cannot meet the obligations, it defaults and liquidates all remaining $z_{i}-x_{i}$ units of assets. Thus, the quantity sold by bank $i$ is $\{p^{-1}(L_{i}-e_{i}-x_{i}Q(\sum_{k=1}^{m}x_{k}))^{+}\}\wedge(z_{i}-x_{i})$ for all $i\in[m]$. Consequently, the market price $p^{\gamma}_{2}$ of illiquid assets at $t=2$ is determined by the following fixed-point equation
$p^{\gamma}_{2}=Q\left(\gamma\sum_{j=1}^{m}x_{j}+\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-x_{j}Q(\sum_{k=1}^{m}x_{k})\right)^{+}}{p^{\gamma}_{2}}\wedge\left(z_{j}-x_{j}\right)\right).$ (4)
Note that a straightforward extension of Lemma 3 in Amini et al. (2016) ensures the uniqueness of $p^{\gamma}_{2}$. We denote each bank $i$’s liquidation size at $t=2$ by
$\psi^{\gamma}_{i}({\bf x})\coloneqq\frac{\left(L_{i}-e_{i}-x_{i}Q(\sum_{k=1}^{m}x_{k})\right)^{+}}{p^{\gamma}_{2}}\wedge\left(z_{i}-x_{i}\right).$ (5)
Finally, bank $i$’s liquid asset value at clearing is given by
$\phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})\coloneqq e_{i}+x_{i}Q\big({\bf 1}^{\top}{\bf x}\big)+\psi^{\gamma}_{i}({\bf x})Q\left(\gamma{\bf 1}^{\top}{\bf x}+\sum_{j=1}^{m}\psi^{\gamma}_{j}({\bf x})\right),$ (6)
where the second and last terms denote bank $i$’s income from fire sales at $t=1$ and $t=2$, respectively.
Remark 1.
Permanent price impact could also be modeled using more sophisticated dynamics, such as, the Volume Weighted Average Price (VWAP) approach investigated in  Banerjee and Feinstein (2021). Unlike static snapshots, this method captures the entire liquidation history by defining the average asset price through an integral of the inverse demand function. However, the technical analysis based on the VWAP framework introduces significant mathematical complexity due to the integration terms; therefore, we retain our current setup to preserve analytical tractability. We anticipate that the qualitative insights derived from our findings would persist under the VWAP framework.
4.2.  Main Problem
Based on the modeling framework described above, we now formulate the decision-making problem for strategic early liquidation. While certain elements of this subsection may overlap with the discussion on robust decision-making in Section 3, we provide here the full technical details omitted there, while extending the formulation to account for permanent price impact.
Robustification against other banks’ actions. As shown in (6), each bank’s liquid asset value at clearing depends on the volume of illiquid assets sold by other banks at $t=1$, which is often unknown. Hence, the solution to the following problem yields bank $i$’s optimal early liquidation decision under the worst-case strategies of other banks:
$\max_{x_{i}}\min_{{\bf x}_{-i}}$ $\phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$ (7)
s.t. ${\bf 0}\leq{\bf x}\leq\bar{{\bf x}},$
where $\bar{x}_{j}$ is defined as in Section 3 and is equal to $\psi^{\gamma}_{j}({\bf 0})$ for any $\gamma\in[0,1]$ and for all $j\in[m]$. The constraint in (7) rules out irrational scenarios where banks liquidate more assets than they would without the early liquidation option, as this would contradict the purpose of the option.
However, solving (7) is extremely difficult because the values $\psi^{\gamma}_{1}({\bf x}),\ldots,\psi^{\gamma}_{m}({\bf x})$ are endogenous to the fixed-point equation in (4), inducing a nested three-level optimization structure in (7). Indeed, one can numerically verify that the problem is highly nonconvex, making it challenging to identify the global optimum even via grid search in low-dimensional settings. To address this issue, we obtain robust upper bounds of $\psi^{\gamma}_{1}({\bf x}),\ldots,\psi^{\gamma}_{m}({\bf x})$:
Lemma 1.
For any ${\bf x}\leq\bar{{\bf x}}$, $\psi^{\gamma}_{j}({\bf x})\leq\bar{x}_{j}-x_{j}$ for all $j\in[m]$ and $\gamma\in[0,1]$.
Several remarks are in order. First, the bound in the preceding lemma is tight and robust against as equality is attained when ${\bf x}={\bf 0}$ or ${\bf x}=\bar{{\bf x}}$. Second, this bound is robust to the early liquidation decisions of other banks; as we shall see shortly, this significantly simplifies the subsequent analysis. Finally, this lemma formalizes the intuition that the early liquidation option does not increase a bank’s total liquidation size.
Relaxation of liquidation size at clearing. Invoking Lemma 1, we circumvent the aforementioned intractability of (7) by relaxing the constraint on each bank $j$’s liquidation size at clearing. Specifically, we allow it to take any value in the interval $0,\bar{x}_{j}-x_{j}]$ rather than restricting it to $\psi_{j}({\bf x})$ for each $j\in[m]$. Under this relaxation, the maximin problem in (7) can be converted into the following form:
$\max_{x_{i},y_{i}}\min_{{\bf x}_{-i},{\bf y}_{-i}}$ $e_{i}+x_{i}Q\big({\bf 1}^{\top}{\bf x}\big)+y_{i}Q\big({\bf 1}^{\top}(\gamma{\bf x}+{\bf y})\big)$ (8)
s.t. ${\bf x}+{\bf y}\leq\bar{{\bf x}}~~\textrm{and}~~{\bf x},{\bf y}\geq{\bf 0}.$
For any feasible solution $({\bf x},{\bf y})$ of (8), Assumption 1 indicates that
$(\bar{x}_{i}-x_{i})Q\big(\gamma x_{i}+\bar{x}_{i}-x_{i}+{\bf 1}^{\top}(\gamma{\bf x}_{-i}+{\bf y}_{-i})\big)$ $\geq y_{i}Q\big(\gamma x_{i}+y_{i}+{\bf 1}^{\top}(\gamma{\bf x}_{-i}+{\bf y}_{-i})\big)$
$\geq y_{i}Q\big(\gamma x_{i}+y_{i}+{\bf 1}^{\top}(\gamma{\bf x}_{-i}+\bar{{\bf x}}_{-i}-{\bf x}_{-i})\big).$
This implies that, first, for fixed ${\bf x}_{-i},{\bf y}_{-i}$, the objective function in (8) is maximized when $y_{i}=\bar{x}_{i}-x_{i}$, and second, for fixed $x_{i},y_{i}$, the objective function is minimized when ${\bf y}_{-i}=\bar{{\bf x}}_{-i}-{\bf x}_{-i}$. In other words, the first constraint of (8) is binding at its optimal solution. Consequently, solving (8) is equivalent to solving
$\max_{x_{i}}\min_{{\bf x}_{-i}}$ $\Phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$ (9)
s.t. ${\bf 0}\leq{\bf x}\leq\bar{{\bf x}},$
where $\Phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})\coloneqq e_{i}+x_{i}Q\big({\bf 1}^{\top}{\bf x}\big)+(\bar{x}_{i}-x_{i})Q\big(\gamma{\bf 1}^{\top}{\bf x}+{\bf 1}^{\top}(\bar{{\bf x}}-{\bf x})\big)$ for all ${\bf x}\in[{\bf 0},\bar{{\bf x}}]$. This is the main problem we aim to investigate in this section, and we refer to the optimal solution to its outer problem as bank $i$’s robust optimal strategy for early liquidation and denote it by $x^{\gamma}_{i}$.
Validity of the formulation. Before presenting the analysis of our problem (9), let us discuss its practical validity to address potential concerns about the implication of optimizing $\Phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$. Specifically, the following result provides insights into how the relaxation affects the liquidation strategy, or equivalently, the relationship between the solutions to (7) and (9)
Proposition 1 (Validity of (9)).
For any fixed $x_{i}\in[0,\bar{x}_{i}]$, the following statements hold:
  1. (i)
    If $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\Phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})\geq L_{i}$, then $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})=L_{i}.$
  2. (ii)
    If $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\Phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})<L_{i}$, then $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\phi^{\gamma}_{i}\left(x_{i},{\bf x}_{-i}\right)\geq\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\Phi^{\gamma}_{i}\left(x_{i},{\bf x}_{-i}\right).$
By the definitions of $\bar{x}_{i}$ and $\psi^{\gamma}_{j}({\bf x})$, it is easy to check that $\phi^{\gamma}_{i}(x_{i},\bar{{\bf x}}_{-i})\leq L_{i}$ for any $x_{i}\in[0,\bar{x}_{i}]$; see the proof of Proposition 1 for more details. [^5] From a practical perspective, this inequality holds since each bank’s liquid asset value at clearing is characterized as either exactly meeting the payment obligation or failing to do so. Thus, the first statement of Proposition 1 shows that if the optimal value of (9) is greater than or equal to $L_{i}$, then bank $i$’s robust optimal strategy, $x^{\gamma}_{i}$, is also an optimal solution to the outer problem of (7). Furthermore, the second statement of the theorem suggests that, even if the optimal value of (9) is smaller than $L_{i}$, bank $i$’s robust optimal strategy still maximizes a lower bound for the worst-case value of its liquid assets at clearing. Accordingly, both statements indicate that the robust optimal strategy obtained by solving (9) can serve as a conservative approximation and a reliable surrogate for the true optimal decision that solves (7). This answers our second research question (Q2), introduced in Section 3, for the case of permanent price impact.
In Figure 1, we numerically validate the results in Proposition 1 by comparing the inner optimal values $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$ and $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\Phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$ across different values of $x_{i}$ in a two-bank system with $m=i=2$ for varying levels of permanent price impact: $\gamma=$ 0, $0.5$, and $1$. The former optimal value is plotted using blue curves—solid for $\gamma=0$, dashed for $\gamma=0.5$, and dotted for $\gamma=1$—without any markers, whereas the latter is plotted using red curves—solid for $\gamma=0$, dashed for $\gamma=0.5$, and dotted for $\gamma=1$—with circle $(\bullet)$ markers. The associated parameter setting can be found in the note of the figure. The left panel of the figure illustrates cases where bank 2’s robust optimal strategy guarantees its solvency regardless of $\gamma$. By contrast, the right panel depicts scenarios in which the proposed strategy leads to solvency when $\gamma=0$ and $0.5$, but default is inevitable when $\gamma=1$.
In both panels, we observe that in the range of $x_{2}$ where the $\bullet$-marked red curves exceed $L_{2}=5$, the unmarked blue curves remain at $L_{2}=5$, which is consistent with Proposition 1(i). Additionally, whenever the $\bullet$-marked red curves are below $L_{2}=5$, they are positioned slightly beneath the unmarked blue curves, which confirms Proposition 1(ii), and the negligible gap between two curves demonstrates the reliability of our relaxed problem (9) at least in this example.
Figure 1: A numerical comparison between the inner optimal values of (7) and (9). We use a two-bank system with Q(x)=10.2xQ(x)=1-0.2x, $L_{1}=L_{2}=5$, $z_{1}=1$, $z_{2}=0.5$, and $e_{2}=4.55$, under $\gamma=0$, $0.5$, and $1$. Also, we set $e_{1}=4.9$ and $e_{1}=4.85$ for the left and right panels, respectively.
4.3.  Analysis of the Robust Optimal Strategy with Permanent Price Impact
In this section, we address our primary research question (Q1) for the case of permanent price impact. To that end, we first derive an analytical expression for the robust optimal strategy $x^{\gamma}_{i}$ and then analyze how it changes with the level of permanent price impact (i.e., $\gamma$).
Derivation of the robust optimal strategy. When $\gamma=0$, (9) reduces to (2), leading to $x^{\gamma}_{i}=x^{*}_{i}=\bar{x}_{i}/2$. In contrast, when $\gamma>0$, early liquidation not only incurs an immediate price impact at $t=1$ but also depresses the asset price at $t=2$. Consequently, from bank $i$’s perspective, the worst-case scenario arises when other banks liquidate more aggressively at $t=1$. To counter this compounded impact at $t=1$, bank $i$ must employ a more prudent early liquidation strategy that depends on the permanency level $\gamma$, resulting in a downward deviation of the robust optimal strategy $x^{\gamma}_{i}$ from $x^{*}_{i}$ that solves (2). This is formalized in the following theorem. Furthermore, the theorem demonstrates that $x^{\gamma}_{i}$ is uniquely determined and admits a (semi-)closed-form expression for any $\gamma\in[0,1]$.
Theorem 1 (Robust Optimal Strategy with Permanent Price Impact).
For each $i\in[m]$ and for each $\gamma\in[0,1]$, the robust optimal strategy $x^{\gamma}_{i}$ that solves (9) is uniquely determined and is bounded from above by $\bar{x}_{i}/2$. More specifically, if $\gamma\leq\gamma^{*}\coloneqq 1-\bar{x}_{i}/(\sum_{j=1}^{n}\bar{x}_{j})$, then $x^{\gamma}_{i}$ is given by
$x^{\gamma}_{i}=\frac{1-\gamma}{2-\gamma}\bar{x}_{i},$ (10)
and if $\gamma>\gamma^{*}$, $x^{\gamma}_{i}$ is the unique solution to the following equation for $x$:
$Q(x+\bar{s}_{i})+xQ^{\prime}(x+\bar{s}_{i})=Q\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)x\big)+(1-\gamma)(\bar{x}_{i}-x)Q^{\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)x\big),$ (11)
where $\bar{s}_{i}=\sum_{j\neq i}\bar{x}_{j}$.
We briefly discuss the technical rationale for the characterization of $x^{\gamma}_{i}$ in the preceding theorem. Analogous to the argument in Section 3, we aggregate the early liquidation sizes of non-target banks into a single quantity, $s_{i}=\sum_{j\neq i}x_{j}$, and rewrite the main problem in (9) as
$\max_{x_{i}\in[0,\bar{x}_{i}]}\min_{s_{i}\in[0,\bar{s}_{i}]}g_{i}^{\gamma}(x_{i},s_{i})=\min_{s_{i}\in[0,\bar{s}_{i}]}\max_{x_{i}\in[0,\bar{x}_{i}]}g_{i}^{\gamma}(x_{i},s_{i}),$ (12)
where $g_{i}^{\gamma}(x,s)\coloneqq e_{i}+xQ(x+s)+(\bar{x}_{i}-x)Q\left(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(x+s)\right)$. The equality in (12) holds due to the concave-convexity of $g_{i}^{\gamma}$ and the minimax theorem (Rockafellar, 2015, Theorem 36.3). Then, the first-order conditions for (12), $({\partial g_{i}^{\gamma}}/{\partial x})(x_{i},s_{i})=0$ and $({\partial g_{i}^{\gamma}}/{\partial s})(x_{i},s_{i})=0$, hold only at
$(x_{i},s_{i})=\left(\frac{1-\gamma}{2-\gamma}\bar{x}_{i},\frac{\gamma\bar{x}_{i}+\bar{s}_{i}}{2-\gamma}\right)$ (13)
and this saddle point is a feasible solution to (12) if and only if $\gamma\leq\gamma^{*}$. This yields the characterization of $x^{\gamma}_{i}$ in (10) for the case where $\gamma\leq\gamma^{*}$. Notably, we observe that (13) satisfies $x_{i}+s_{i}=\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(x_{i}+s_{i})$. This implies that (13) effectively balances the resulting price impact at $t=1$ and $t=2$, as the left- and right-hand sides of this identity correspond to the arguments of the function $Q$ in the definition of $g_{i}^{\gamma}$.
Conversely, when $\gamma>\gamma^{*}$, one can easily intuit that the optimal solution to (12) lies at the boundary $s_{i}=\bar{s}_{i}$ because the unique saddle point (13) falls within $0,\bar{x}_{i}/2)\times(\bar{s}_{i},\bar{x}_{i}+\bar{s}_{i}]$. This implies that if the permanent price impact is sufficiently large, the worst-case scenario for bank $i$ arises when other banks liquidate their entire required volume of illiquid assets only at $t=1$. In this case, the robust optimal strategy $x^{\gamma}_{i}$ is characterized by the following first-order condition $({\partial g_{i}^{\gamma}}/{\partial x})(x_{i},\bar{s}_{i})=0$, which is equivalent to solving the equation in (11).
While the reasoning outlined above provides a straightforward path to constructing a solution to (9), the essence of Theorem 1 lies in establishing its uniqueness, thereby ruling out the possibility of any alternative solutions; see the proof of the theorem in Section A for more details. It is also worth highlighting that solving (11) is computationally tractable since $(\partial g_{i}^{\gamma}/\partial x)(\cdot,\bar{s}_{i})$ is strictly decreasing. Moreover, when the inverse demand function is linear—a common assumption in prior studies—the solution admits a closed form. Specifically, if $Q(x)=P-\alpha x$, then (11) can be written as $P-\alpha(2x+\bar{s}_{i})=P-\alpha((2-\gamma)\bar{x}_{i}+\gamma\bar{s}_{i}-2(1-\gamma)x)$, and thus,
$x^{\gamma}_{i}=\frac{\bar{x}_{i}}{2}-\frac{(1-\gamma)\bar{s}_{i}}{2(2-\gamma)}.$ (14)
Accordingly, in this linear setting, $x^{\gamma}_{i}$ exhibits a V-shaped dependence on $\gamma$: it decreases in $\gamma$ on $0,\gamma^{*})$, as shown in (10), and increases with $\gamma$ on $(\gamma^{*},1]$ as indicated by (14).
Structural dependence on the parameter $\gamma$. The V-shaped dependence of the robust optimal strategy mentioned above is not limited to the specific choice of $Q(\cdot)$. Indeed, the following theorem demonstrates that, under a mild condition, this behavior holds for “all” inverse demand functions satisfying Assumption 1.
Theorem 2 (Impact of $\gamma$: A Typical Scenario).
For any $i\in[m]$, if $\gamma^{*}\geq 1/2$, then the robust optimal strategy $x^{\gamma}_{i}$ increases in $\gamma$ on $(\gamma^{*},1]$.
Firstly, we note that the condition $\gamma^{*}\geq 1/2$ is equivalent to $\bar{x}_{i}\leq\bar{s}_{i}\coloneqq\sum_{j\neq i}\bar{x}_{j}$, meaning that bank $i$’s contribution to the price impact is smaller than the aggregate contribution of all other banks. Thus, it is typically satisfied in practice. Secondly, as discussed earlier, the worst-case scenario for bank $i$ with $\gamma>\gamma^{*}$ corresponds to a situation in which all other banks collectively liquidate $\bar{s}_{i}$ units of illiquid assets exclusively at $t=1$. Therefore, Theorem 2 aligns with the intuition that, given $\bar{x}_{i}\leq\bar{s}_{i}$, a larger $\gamma$ leads to a sharper price decline at $t=2$. This prompts bank $i$ to increase sales at $t=1$ to avoid the anticipated future price distress. By combining this result with the analytical solution (10) for $x^{\gamma}_{i}$ when $\gamma\leq\gamma^{*}$, the aforementioned V-shaped dependence holds generally.
The left and right panels of Figure 2 depict a bank’s robust optimal strategy and its worst-case scenario on the aggregate early liquidation size of other banks, respectively, as functions of $\gamma$ when the inverse demand function $Q$ is nonlinear; specifically $Q(x)=\exp(-0.2x)$. In this example, $\gamma^{*}\approx 0.7$, satisfying the condition in Theorem 2. Consistent with our theoretical results, the left panel displays the aforementioned non-monotonic behavior of $x^{\gamma}_{i}$ with the tipping point $\gamma^{*}$ and the upper bound $\bar{x}_{i}/2$, and the right panel shows that under the worst-case scenario, higher $\gamma$ drives other banks to liquidate more at $t=1$ when $\gamma\leq\gamma^{*}$, and to liquidate exclusively in the first period once $\gamma>\gamma^{*}$.
Figure 2: A bank’s robust optimal strategy and its worst-case scenario on the aggregate early liquidation size of other banks under various levels of permanent price impact when $\gamma^{*}$ is large. We use a four-bank system with $i=1$, $L_{1}=\cdots=L_{4}=6$, $z_{1}=\cdots=z_{4}=1$, $w_{1}=0.3$, $w_{2}=w_{3}=0.5$, and $w_{4}=0.8$. We set $Q(x)=\exp(-0.2x)$ and $s(x)=\mathop{\arg\min}_{s\in[0,\bar{s}_{i}]}g_{i}^{\gamma}(x,s)$, where $g_{i}^{\gamma}$ is defined as in (12).
However, one can easily intuit that the said V-shaped dependence does not hold universally. Particularly, in an extreme case where bank $i$’s contribution to the price impact is significantly greater than the aggregate contribution of all other banks (i.e., $\bar{x}_{i}\gg\bar{s}_{i}$), the price impact would be predominantly driven by bank $i$’s own liquidation. Hence, as $\gamma$ increases, the optimal response for bank $i$ would be to monotonically reduce its early liquidation size to alleviate the effects of persistent price impact. This intuition is rigorously verified in the following proposition.
Proposition 2 (Impact of $x^{\gamma}_{i}$: An Extreme Scenario).
For any $i\in[m]$, if $Q^{\prime\prime}(\cdot)>0$ and $\gamma^{*}$ is sufficiently small, then the robust optimal strategy $x^{\gamma}_{i}$ decreases in $\gamma$ on $[0,1]$.
In Figure 3, we conduct an experiment similar to that in Figure 2, but with $\gamma^{*}\approx 0.11$. In this case, the non-monotonicity vanishes; instead, $x^{\gamma}_{i}$ decreases monotonically with $\gamma$, validating Theorem 2. The right panel of the figure mirrors the pattern in that of Figure 2: increasing with $\gamma$ until $\gamma\leq\gamma^{*}$ and plateauing at $\bar{s}_{i}$ thereafter.
Figure 3: A bank’s robust optimal strategy and its worst-case scenario on the aggregate early liquidation size of other banks under various levels of permanent price impact when $\gamma^{*}$ is small. We use a four-bank system with $i=1$, $L_{1}=\cdots=L_{4}=6$, $z_{1}=4$, $z_{2}=z_{3}=z_{4}=0.6$, $w_{1}=0.3$, $w_{2}=w_{3}=0.5$, and $w_{4}=0.8$. We set $Q(x)=\exp(-0.2x)$ and $s(x)=\mathop{\arg\min}_{s\in[0,\bar{s}_{i}]}g_{i}^{\gamma}(x,s)$, where $g_{i}^{\gamma}$ is defined as in (12).
5.  Robust Optimal Strategy with Interbank Exposures
We now take into account debt exposures between banks and investigate how they affect the robust optimal strategy. Below, we present the associated model, which extends the framework introduced in Cifuentes et al. (2005) and used in many follow-up studies (Chen et al., 2016; Amini et al., 2016; Weber and Weske, 2017; Bernard et al., 2022). Our extension accounts for (i) the possibility of early liquidation and (ii) each bank’s liabilities to external entities beyond the financial system.
Table 2 describes the initial balance sheet of each bank $i$ in an $m$-bank system operating over two periods. The difference from Table 1 lies in that interbank assets $\{L_{ji}\}_{j\neq i}$ are added to the asset side, while total liabilities $L_{i}$ are now composed of liabilities to external entities (denoted by $b_{i}>0$) and interbank liabilities $\{L_{ij}\}_{j\neq i}$, i.e., $L_{i}=b_{i}+\sum_{j\neq i}L_{ij}$. Here, $L_{ij}$ denotes the nominal liabilities of bank $i$ owing to bank $j$, with $L_{ii}=0$, for all $i,j\in[m]$. We again assume that all balance sheet components are assumed to remain unchanged over the two periods ($t=1,2$) and all liabilities are cleared at $t=2$.
Table 2: Initial balance sheet of bank $i$ with interbank exposures before liquidation
Assets Liabilities and Net Worth
Liquid assets $e_{i}$ External liabilities $b_{i}$
Interbank assets $L_{ji},j\neq i$ Interbank liabilities $L_{ij},j\neq i$
Illiquid assets $z_{i}$ with price $P$ Net worth $e_{i}+\sum_{j\neq i}L_{ji}+z_{i}P-L_{i}$
The component $L_{i}$ represents bank $i$’s total liabilities, given by $L_{i}=b_{i}+\sum_{j\neq i}L_{ij}$.
It is also assumed that all liabilities are of equal seniority; if a bank’s asset value at clearing is insufficient to cover total liabilities, its remaining assets are distributed to creditors on a pro rata basis. To illustrate this, we introduce the relative liability matrix ${\bf A}=(A_{ij})$, where $A_{ij}=L_{ij}/L_{i}$ represents the proportion of bank $i$’s liabilities to bank $j$. Then, if bank $j$’s total debt repayment is $\pi_{j}$ for all $j\in[m]$, the total payment received by bank $i$ (i.e., the value of its interbank assets) becomes $\sum_{j=1}^{m}A_{ji}\pi_{j}$ (or equivalently, ${\bf A}_{i}^{\top}\boldsymbol{\pi}$). Consequently, given that each bank $i$ sells off $x_{i}$ units of illiquid assets at $t=1$, the market price $p_{2}^{\bf A}$ of illiquid assets at $t=2$ and each bank $i$’s clearing payment $\pi_{j}^{\bf A}$ are defined as the solution to the fixed-point equations, using the inverse demand function $Q$ satisfying Assumption 1 over two periods: [^6] In this section, we set aside permanent price impact in order to focus exclusively on the impact of interbank exposures.
$\left\{~\begin{aligned} &p_{2}^{\bf A}=Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-x_{j}Q\big({\bf 1}^{\top}{\bf x}\big)-{\bf A}_{j}^{\top}\boldsymbol{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)\right),\\ &\pi_{i}^{\bf A}=L_{i}\wedge\left(e_{i}+x_{i}Q\big({\bf 1}^{\top}{\bf x}\big)+(z_{i}-x_{i})p_{2}^{\bf A}+{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}\right),~~i=1,\ldots,m.\end{aligned}\right$. (15)
Similar to Section 4, the existence and uniqueness of the above solution is guaranteed by Amini et al. (2016). Each bank $i$’s liquidation size at $t=2$ is then given by
$\psi_{i}^{\bf A}({\bf x})\coloneqq\frac{\left(L_{i}-e_{i}-x_{i}Q({\bf 1}^{\top}{\bf x})-{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{i}-x_{i}\right).$ (16)
Finally, bank $i$’s liquid asset value at clearing is defined as
$\phi^{\bf A}_{i}(x_{i},{\bf x}_{-i})\coloneqq e_{i}+x_{i}Q\big({\bf 1}^{\top}{\bf x}\big)+\psi_{i}^{\bf A}({\bf x})Q\big({\bf 1}^{\top}\boldsymbol{\psi}^{\bf A}({\bf x})\big)+{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}.$ (17)
The subsequent analyses proceed in parallel with Section 4; we (i) relax the liquidation size at clearing, (ii) formulate the robust optimization problem (i.e., the counterpart of (9)), (iii) verify the validity of this formulation, and (iv) derive the robust optimal strategy, which is the optimal solution to this problem. Although the theoretical results here mirror those in Section 4, the proofs are nontrivial extensions and differ significantly due to the presence of $\boldsymbol{\pi}^{\bf A}$, which is characterized by a fixed-point equation.
Let us denote $\bar{x}^{\bf A}_{i}\coloneqq\psi_{i}^{\bf A}({\bf 0})$ for each $i\in[m]$, which represents bank $i$’s liquidation size at clearing in the absence of early liquidation, and we assume $\bar{{\bf x}}^{\bf A}>{\bf 0}$ since otherwise liquidation is not required for some banks. We then establish a variant of Lemma 1 that gives a robust bound of the liquidation size at $t=2$:
Lemma 2.
For any ${\bf x}\leq\bar{{\bf x}}^{\bf A}$, $\psi_{j}^{\bf A}({\bf x})\leq\bar{x}^{\bf A}_{j}-x_{j}$ for all $j\in[m]$.
Based on the above lemma, given a realization ${\bf x}\leq\bar{{\bf x}}^{\bf A}$ of the early liquidation size, we can identify each bank $i$’s maximum possible income from fire sales at $t=1$ and $t=2$ as follows:
$c_{i}({\bf x})\coloneqq x_{i}Q\big({\bf 1}^{\top}{\bf x}\big)+\big(\bar{x}^{\bf A}_{i}-x_{i}\big)Q\Big({\bf 1}^{\top}\big(\bar{{\bf x}}^{\bf A}-{\bf x}\big)\Big).$ (18)
This provides an upper bound on the sum of the middle two terms on the right-hand side of (17), reflecting the worst-case scenario in which banks ultimately liquidate the same volume of illiquid assets as they would without early liquidation. Then, analogous to the construction of the second equation in (15), we can characterize the vector of clearing payments associated with this worst-case scenario, denoted by $\boldsymbol{\pi}({\bf c}({\bf x}))$, as the solution to the following fixed-point equation:
$\boldsymbol{\pi}={\bf L}\wedge\left({\bf e}+{\bf c}({\bf x})+{\bf A}^{\top}\boldsymbol{\pi}\right),$ (19)
where ${\bf c}(\cdot)\coloneqq\big(c_{1}(\cdot),\dots,c_{m}(\cdot)\big)^{\top}$ and $\boldsymbol{\pi}({\bf c}({\bf x}))$ is unique for any ${\bf x}\in[{\bf 0},\bar{{\bf x}}^{\bf A}]$ by Theorem 2 of Eisenberg and Noe (2001).
Consequently, the counterpart of (9) for the case with interbank exposures can be formulated as follows:
$\max_{x_{i}}\min_{{\bf x}_{-i}}$ $\Phi^{\bf A}_{i}(x_{i},{\bf x}_{-i})$ (20)
s.t. ${\bf 0}\leq{\bf x}\leq\bar{{\bf x}}^{\bf A},$
where $\Phi^{\bf A}_{i}(x_{i},{\bf x}_{-i})\coloneqq e_{i}+c_{i}({\bf x})+{\bf A}_{i}^{\top}\boldsymbol{\pi}\big({\bf c}({\bf x})\big)$ consists of the initial external liquid assets $e_{i}$, the maximum possible liquidation income $c_{i}({\bf x})$, and the worst-case interbank assets ${\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf c}({\bf x}))$. In line with Proposition 1, we demonstrate the validity of the above formulation in the following proposition, which addresses (Q2) in Section 3 for the case of interbank exposures:
Proposition 3 (Validity of (20)).
For any fixed $x_{i}\in[0,\bar{x}_{i}]$, the following statements hold:
  1. (i)
    If $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(x_{i},{\bf x}_{-i})\geq L_{i}$, then $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\phi^{\bf A}_{i}(x_{i},{\bf x}_{-i})=L_{i}.$
  2. (ii)
    If $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(x_{i},{\bf x}_{-i})<L_{i}$, then $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\phi^{\bf A}_{i}\left(x_{i},{\bf x}_{-i}\right)\geq\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}\left(x_{i},{\bf x}_{-i}\right).$
A salient feature of the problem involving interbank exposures—distinguishing it from the previous case—is that each bank must not only avoid extreme price impacts on its own assets over the two periods but also seek to mitigate price impacts on other banks to increase the payments it receives from them. Furthermore, these payments are implicitly characterized by the fixed-point equation (19), making the objective function in (20) neither convex nor concave in $x_{i}$ or ${\bf x}_{-i}$. This precludes the use of the minimax theorem and the approach of aggregating other banks’ early liquidation sizes, significantly complicating the technical analysis required to identify the optimal solution to (20). Nevertheless, the following theorem successfully characterizes the robust optimal strategy for this problem, which turns out to take a surprisingly simple form, structurally analogous to that of the illustrative example in Section 3.
Theorem 3 (Robust Optimal Strategy with Interbank Exposures).
For each $i\in[m]$, the robust optimal strategy $x_{i}^{\bf A}$ that solves (20) is uniquely determined as $x_{i}^{\bf A}=\bar{x}^{\bf A}_{i}/2$.
Despite the aforementioned technical complications, the simple characterization of the robust optimal strategy is attributed to our finding that the sensitivity of $\boldsymbol{\pi}({\bf c}({\bf x}))$ to $\bf x$} is dominated by that of $c_{i}({\bf x})$ under bank $i$’s worst-case scenario. Consequently, the robust optimal strategy is primarily driven by a solution to the problem $\max_{x_{i}\in[0,\bar{x}^{\bf A}_{i}]}\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}c_{i}({\bf x})$. By applying symmetry and the first-order conditions, its solution is found to be $(\bar{x}^{\bf A}_{i}/2,\bar{{\bf x}}^{\bf A}_{-i}/2)$, leading to the above result.
Practically speaking, if bank $i$ were to adopt an overly aggressive or conservative strategy for its early liquidation, the associated price impact under the worst-case scenario would generate negative externalities across the entire financial system, leading to significant losses in counterparties’ liquidation income. Although a minority of these institutions may derive limited gains through asset fire sales, the resulting repayments from this group would be insufficient to counterbalance the aggregate shortfall. Hence, as bank $i$ would fail to hedge its own liquidations costs through incoming interbank payments, the robust optimal strategy converges toward a moderate and balanced early liquidation solution: selling half of the assets that would otherwise be liquidated under the no-early-liquidation benchmark. This finally addresses (Q1) in Section 3 for the case involving interbank exposures.
6.  Conclusion
In this paper, we provide new insights into an individual bank’s early liquidation strategy in the context of systemic risk. Using a stylized two-period model, we formulate a maximin problem that maximizes the worst-case value of the bank’s liquid assets at clearing, accounting for the uncertain liquidation decisions of counterparties. To tackle the analytical intractability of the said problem, we propose a relaxed version that is both practically valid and analytically tractable. Crucially, we show that this relaxation yields a unique, robust optimal strategy that admits a (semi-)closed-form expression. We further highlight the distinct effects of two key factors widely investigated in the systemic risk literature: permanent price impact and interbank exposures. Specifically, permanent price impact leads to structural changes in the robust optimal strategy, which typically exhibits a non-monotonic (V-shaped) dependence on the magnitude of the impact. In contrast, interbank exposures play a less significant role and do not drive such structural shifts.
Several directions remain open for future research. Firstly, while our robust optimal strategy is explicitly characterized and offers insights into early liquidation based on the relaxed maximin problem, it remains inconclusive whether a tighter yet analytically tractable relaxation exists. Addressing this question may require more precise and explicit approximations of the dependence among liquidation sizes across two periods. Secondly, although our framework accommodates permanent price impact and interbank liabilities individually, integrating both features into a unified framework poses substantial analytical challenges such as nonconvexity and asymmetry. These issues prevent a direct extension of our current analysis and hence represent a promising path for subsequent studies. Lastly, we develop the two-period liquidation strategy without delving into the quantification of other sources of financing risks. In practice, institutions generally evaluate various risk factors comprehensively. For example, apart from liquidation, illiquid assets can also be used in collateralized borrowing. Therefore, further investigation of a versatile financing strategy would be an interesting research direction.
Appendix A Proofs of Theoretical Results in Section 4
We first introduce several inequalities that will be useful in the proofs of our theoretical results. Specifically, for all $x>0,y\geq 0$ with $x+y<z_{tot}$, let $f(x,y)=xQ(x+y)$ where $Q$ is the inverse demand function. Then, under Assumption 1, it is straightforward to check that
$f_{x}(x,y)\coloneqq\frac{\partial f(x,y)}{\partial x}=Q(x+y)+xQ^{\prime}(x+y)>0,$ (21a)
$f_{y}(x,y)\coloneqq\frac{\partial f(x,y)}{\partial y}=xQ^{\prime}(x+y)<0,$ (21b)
$f_{xx}(x,y)\coloneqq\frac{\partial^{2}f(x,y)}{\partial x^{2}}=2Q^{\prime}(x+y)+xQ^{\prime\prime}(x+y)<0,$ (21c)
$f_{xy}(x,y)\coloneqq\frac{\partial^{2}f(x,y)}{\partial x\partial y}=Q^{\prime}(x+y)+xQ^{\prime\prime}(x+y)<0.$ (21d)
Proof.
Proof of Lemma 1. Let $\bar{p}\coloneqq Q({\bf 1}^{\top}\bar{{\bf x}})$. By (4) and (5), it is easy to check that
$\bar{x}_{j}\bar{p}=\psi^{\gamma}_{j}({\bf 0})Q\left({\bf 1}^{\top}\psi^{\gamma}_{j}({\bf 0})\right)=(L_{j}-e_{j})\wedge(z_{j}\bar{p})\leq L_{j}-e_{j}.$ (22)
Assume by contradiction that $\bar{p}>p^{\gamma}_{2}$, where $p^{\gamma}_{2}$ is defined as in (4). Then, we first observe that $p_{1}\coloneqq Q({\bf 1}^{\top}{\bf x})\geq Q({\bf 1}^{\top}\bar{{\bf x}})=\bar{p}>p^{\gamma}_{2}$ since $Q(\cdot)$ is strictly decreasing by Assumption 1(ii). This implies that for any $\gamma\in[0,1]$,
$\left(L_{j}-e_{j}-x_{j}p_{1}\right)^{+}+\gamma x_{j}p^{\gamma}_{2}=\left[L_{j}-e_{j}-x_{j}\left(p_{1}-\gamma p^{\gamma}_{2}\right)\right]\vee\left(\gamma x_{j}p^{\gamma}_{2}\right)\leq(L_{j}-e_{j})\vee(x_{j}\bar{p})=L_{j}-e_{j},$ (23)
where the last equality holds by (22) and $u\vee v\coloneqq\max\{u,v\}$ for any $u,v\in\mathbb{R}$.
Let
$\rho_{0}=Q\left(\sum_{j=1}^{m}\frac{L_{j}-e_{j}}{p^{\gamma}_{2}}\wedge z_{j}\right)~~\text{and}~~\rho_{k+1}=Q\left(\sum_{j=1}^{m}\frac{L_{j}-e_{j}}{\rho_{k}}\wedge z_{j}\right)~\text{for}~k=0,1,2,\ldots$ (24)
Then, by (23) and Assumption 1(ii), we have
$p^{\gamma}_{2}$ $=Q\left(\gamma\sum_{j=1}^{m}x_{j}+\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-x_{j}p_{1}\right)^{+}}{p^{\gamma}_{2}}\wedge(z_{j}-x_{j})\right)$
$=Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-x_{j}p_{1}\right)^{+}+\gamma x_{j}p^{\gamma}_{2}}{p^{\gamma}_{2}}\wedge\big(z_{j}-(1-\gamma)x_{j}\big)\right)$
$\geq Q\left(\sum_{j=1}^{m}\frac{L_{j}-e_{j}}{p^{\gamma}_{2}}\wedge z_{j}\right)$
$=\rho_{0}.$
Furthermore, if $\rho_{k}\leq p^{\gamma}_{2}$ for some $k\geq 0$, then
$\rho_{k+1}=Q\left(\sum_{j=1}^{m}\frac{L_{j}-e_{j}}{\rho_{k}}\wedge z_{j}\right)\leq Q\left(\sum_{j=1}^{m}\frac{L_{j}-e_{j}}{p^{\gamma}_{2}}\wedge z_{j}\right)=p^{\gamma}_{2}.$ (25)
Hence, by induction, we obtain that $\rho_{k}\leq p^{\gamma}_{2}$ for all $k$, which suggests that $\bar{p}\leq p^{\gamma}_{2}$ since $\rho_{k}\rightarrow\bar{p}$ as $k$ increases. This contradicts the assumption and indicates that $\bar{p}\leq p^{\gamma}_{2}$.
Consequently, the following holds for any $\gamma\in[0,1]$:
$x_{j}+\psi^{\gamma}_{j}({\bf x})$ $=x_{j}+\frac{\left(L_{j}-e_{j}-x_{j}p_{1}\right)^{+}}{p^{\gamma}_{2}}\wedge\left(z_{j}-x_{j}\right)$
$\leq x_{j}+\frac{\left(L_{j}-e_{j}-x_{j}p_{1}\right)^{+}}{\bar{p}}\wedge\left(z_{j}-x_{j}\right)$
$=\frac{\big(L_{j}-e_{j}-x_{j}(p_{1}-\bar{p})\big)\vee\left(x_{j}\bar{p}\right)}{\bar{p}}\wedge z_{j}$
$\leq\frac{\left(L_{j}-e_{j}\right)\vee\left(\bar{x}_{j}\bar{p}\right)}{\bar{p}}\wedge z_{j}$
$=\frac{L_{j}-e_{j}}{\bar{p}}\wedge z_{j}$
$=\bar{x}_{j},$
which completes the proof. ∎
Proof.
Proof of Proposition 1. Fix $x_{i}\leq\bar{x}_{i}$ and let ${\bf u}^{\gamma}\coloneqq\mathop{\arg\min}_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$. By the definition of $\bar{{\bf x}}$ and (21a), we observe that $e_{i}+x_{i}Q(x_{i}+{\bf 1}^{\top}\bar{{\bf x}}_{-i})\leq e_{i}+\bar{x}_{i}Q({\bf 1}^{\top}\bar{{\bf x}})\leq L_{i}$. Thus, we have
$\phi^{\gamma}_{i}(x_{i},{\bf u}^{\gamma})$ $\leq\phi^{\gamma}_{i}(x_{i},\bar{{\bf x}}_{-i})$ (26)
$=e_{i}+x_{i}Q\big(x_{i}+{\bf 1}^{\top}\bar{{\bf x}}_{-i}\big)+\left(L_{i}-e_{i}-x_{i}Q\big(x_{i}+{\bf 1}^{\top}\bar{{\bf x}}_{-i}\big)\right)\wedge\left((z_{i}-x_{i})p^{\gamma}_{2}\right)$
$\leq e_{i}+x_{i}Q\big(x_{i}+{\bf 1}^{\top}\bar{{\bf x}}_{-i}\big)+L_{i}-e_{i}-x_{i}Q\big(x_{i}+{\bf 1}^{\top}\bar{{\bf x}}_{-i}\big)$
$=L_{i},$
where the first equality holds by (5) to (6).
We now assume $\phi^{\gamma}_{i}(x_{i},{\bf u}^{\gamma})<L_{i}$ and define ${\bf v}^{\gamma}\in\mathbb{R}^{m}$ with $v_{i}^{\gamma}=x_{i}$ and ${\bf v}_{-i}^{\gamma}={\bf u}^{\gamma}$. Then, by (6), it follows that
$\psi^{\gamma}_{i}({\bf v}^{\gamma})=\frac{\phi^{\gamma}_{i}(x_{i},{\bf u}^{\gamma})-e_{i}-x_{i}Q\left({\bf 1}^{\top}{\bf v}^{\gamma}\right)}{Q\big(\gamma{\bf 1}^{\top}{\bf v}^{\gamma}+\sum_{j=1}^{m}\psi^{\gamma}_{j}({\bf v}^{\gamma})\big)}<\frac{\big(L_{i}-e_{i}-x_{i}Q({\bf 1}^{\top}{\bf v}^{\gamma})\big)^{+}}{Q\big(\gamma{\bf 1}^{\top}{\bf v}^{\gamma}+\sum_{j=1}^{m}\psi^{\gamma}_{j}({\bf v}^{\gamma})\big)}.$ (27)
Hence, $\psi^{\gamma}_{i}({\bf v}^{\gamma})=z_{i}-x_{i}$ by (5). Since $\bar{x}_{i}\leq z_{i}$ by the definition of $\bar{x}_{i}$ and $\psi^{\gamma}_{j}({\bf v}^{\gamma})\leq\bar{x}_{j}-x_{j}$ for all $j$ by Lemma 1, we obtain
$\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\Phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})$ $\leq\Phi^{\gamma}_{i}\left(x_{i},{\bf u}^{\gamma}\right)$ (28)
$=e_{i}+x_{i}Q\big({\bf 1}^{\top}{\bf v}^{\gamma}\big)+\left(\bar{x}_{i}-x_{i}\right)Q\Big(\gamma{\bf 1}^{\top}{\bf v}^{\gamma}+{\bf 1}^{\top}\big(\bar{{\bf x}}-{\bf v}^{\gamma}\big)\Big)$
$\leq e_{i}+x_{i}Q\big({\bf 1}^{\top}{\bf v}^{\gamma}\big)+(z_{i}-x_{i})Q\Bigg(\gamma{\bf 1}^{\top}{\bf v}^{\gamma}+\sum_{j=1}^{m}\psi^{\gamma}_{j}({\bf v}^{\gamma})\Bigg)$
$=e_{i}+x_{i}Q\big({\bf 1}^{\top}{\bf v}^{\gamma}\big)+\psi^{\gamma}_{i}({\bf v}^{\gamma})Q\Bigg(\gamma{\bf 1}^{\top}{\bf v}^{\gamma}+\sum_{j=1}^{m}\psi^{\gamma}_{j}({\bf v}^{\gamma})\Bigg)$
$=\phi^{\gamma}_{i}(x_{i},{\bf u}^{\gamma}).$
Consequently, $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\Phi^{\gamma}_{i}(x_{i},{\bf x}_{-i})<L_{i}$, which demonstrates the first statement by contrapositive.
The second statement is straightforward if $\phi^{\gamma}_{i}(x_{i},{\bf u}^{\gamma})=L_{i}$. Otherwise, it holds by (28). ∎
Proof.
Proof of Theorem 1. Fix $i\in[m]$. For any $x\in[0,\bar{x}_{i}]$ and $s\in[0,\bar{s}_{i}]$, we define
$g_{i}^{\gamma}(x,s)\coloneqq e_{i}+xQ(x+s)+(\bar{x}_{i}-x)Q(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(x+s)).$
Then, (9) can be rewritten as
$\max_{x\in[0,\,\bar{x}_{i}]}\,\min_{s\in\left[0,\,\bar{s}_{i}\right]}\;g_{i}^{\gamma}(x,\,s).$ (29)
Using this reformulation, we prove the theorem in three steps.
Step 1. Let $\tilde{x}_{i}^{\gamma}$ be the value satisfying (11) with respect to $x$. We establish its existence and uniqueness given $\gamma\geq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$. By Assumption 1(iv) and (21c), we observe that $g_{i}^{\gamma}(\cdot,\bar{s}_{i})$ is strictly concave on $0,\bar{x}_{i}]$. Also, it is easy to see that (11) holds if and only if
$\frac{\partial g_{i}^{\gamma}}{\partial x}\left(\tilde{x}_{i}^{\gamma},\bar{s}_{i}\right)=0.$ (30)
Furthermore, a straightforward calculation yields that
$\frac{\partial g_{i}^{\gamma}}{\partial x}\left(\frac{(1-\gamma)\bar{x}_{i}}{2-\gamma},\bar{s}_{i}\right)$ $=Q\left(\frac{(1-\gamma)\bar{x}_{i}}{2-\gamma}+\bar{s}_{i}\right)+\frac{(1-\gamma)\bar{x}_{i}}{2-\gamma}Q^{\prime}\left(\frac{(1-\gamma)\bar{x}_{i}}{2-\gamma}+\bar{s}_{i}\right)$ (31)
$~~~~-Q\left(\frac{(1-\gamma)\bar{x}_{i}}{2-\gamma}+\gamma(\bar{x}_{i}+\bar{s}_{i})\right)-\frac{(1-\gamma)\bar{x}_{i}}{2-\gamma}Q^{\prime}\left(\frac{(1-\gamma)\bar{x}_{i}}{2-\gamma}+\gamma(\bar{x}_{i}+\bar{s}_{i})\right),$
and thus, by (21d), this value is non-negative if and only if $\gamma\geq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$. Similarly, if $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$, by Assumption 1 (iv), we obtain
$\frac{\partial g_{i}^{\gamma}}{\partial x}\left(\frac{\bar{x}_{i}}{2}-\frac{(1-\gamma)\bar{s}_{i}}{2(2-\gamma)},\bar{s}_{i}\right)$ (32)
$=Q\left(\frac{\bar{x}_{i}}{2}+\frac{(3-\gamma)\bar{s}_{i}}{2(2-\gamma)}\right)+\left(\frac{\bar{x}_{i}}{2}-\frac{(1-\gamma)\bar{s}_{i}}{2(2-\gamma)}\right)Q^{\prime}\left(\frac{\bar{x}_{i}}{2}+\frac{(3-\gamma)\bar{s}_{i}}{2(2-\gamma)}\right)-Q\left(\frac{(1+\gamma)\bar{x}_{i}}{2}+\frac{(1+2\gamma-\gamma^{2})\bar{s}_{i}}{2(2-\gamma)}\right)$
$~~~~-\frac{(1-\gamma)(2-\gamma)\bar{x}_{i}+(1-\gamma)^{2}\bar{s}_{i}}{2(2-\gamma)}Q^{\prime}\left(\frac{(1+\gamma)\bar{x}_{i}}{2}+\frac{(1+2\gamma-\gamma^{2})\bar{s}_{i}}{2(2-\gamma)}\right)$
$\leq\left(\frac{\bar{x}_{i}}{2}-\frac{(1-\gamma)\bar{s}_{i}}{2(2-\gamma)}\right)Q^{\prime}\left(\frac{\bar{x}_{i}}{2}+\frac{(3-\gamma)\bar{s}_{i}}{2(2-\gamma)}\right)-\frac{\gamma\bar{x}_{i}-(1-\gamma)\bar{s}_{i}}{2}Q^{\prime}\left(\frac{\bar{x}_{i}}{2}+\frac{(3-\gamma)\bar{s}_{i}}{2(2-\gamma)}\right)$
$~~~~-\frac{(1-\gamma)(2-\gamma)\bar{x}_{i}+(1-\gamma)^{2}\bar{s}_{i}}{2(2-\gamma)}Q^{\prime}\left(\frac{(1+\gamma)\bar{x}_{i}}{2}+\frac{(1+2\gamma-\gamma^{2})\bar{s}_{i}}{2(2-\gamma)}\right)$
$\leq\left(\frac{\bar{x}_{i}}{2}-\frac{(1-\gamma)\bar{s}_{i}}{2(2-\gamma)}\right)Q^{\prime}\left(\frac{\bar{x}_{i}}{2}+\frac{(3-\gamma)\bar{s}_{i}}{2(2-\gamma)}\right)-\frac{\gamma\bar{x}_{i}-(1-\gamma)\bar{s}_{i}}{2}Q^{\prime}\left(\frac{\bar{x}_{i}}{2}+\frac{(3-\gamma)\bar{s}_{i}}{2(2-\gamma)}\right)$
$~~~~-\frac{(1-\gamma)(2-\gamma)\bar{x}_{i}+(1-\gamma)^{2}\bar{s}_{i}}{2(2-\gamma)}Q^{\prime}\left(\frac{\bar{x}_{i}}{2}+\frac{(3-\gamma)\bar{s}_{i}}{2(2-\gamma)}\right)$
$=0,$
where the second inequality holds since
$\frac{\bar{x}_{i}}{2}+\frac{(3-\gamma)\bar{s}_{i}}{2(2-\gamma)}<\frac{(1+\gamma)\bar{x}_{i}}{2}+\frac{(1+2\gamma-\gamma^{2})\bar{s}_{i}}{2(2-\gamma)}~~\text{for all}~\gamma\geq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i}).$
By combining all the above results and the continuity of $\partial g_{i}^{\gamma}/\partial x(\cdot,\bar{s}_{i})$, $\tilde{x}_{i}^{\gamma}$ exists and is unique if $\gamma\geq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$, and we have
$\frac{(1-\gamma)\bar{x}_{i}}{2-\gamma}\leq\tilde{x}_{i}^{\gamma}\leq\frac{\bar{x}_{i}}{2}-\frac{(1-\gamma)\bar{s}_{i}}{2(2-\gamma)}~~\text{if}~~\gamma\geq\frac{\bar{s}_{i}}{\bar{x}_{i}+\bar{s}_{i}}.$ (33)
Step 2. Let
$z^{\gamma}\coloneqq\left\{\begin{aligned} &\frac{(1-\gamma)\bar{x}_{i}}{2-\gamma}&&~\text{if}~\gamma\leq\frac{\bar{s}_{i}}{\bar{x}_{i}+\bar{s}_{i}};\\ &\tilde{x}_{i}^{\gamma}&&~\text{otherwise.}\end{aligned}\right$.
For each $x\in[0,\bar{x}_{i}]$, define ${\cal S}(x)=\mathop{\arg\min}\limits_{s\in[0,\bar{s}_{i}]}g_{i}^{\gamma}(x,s)$ and
$s(x)=\left\{\begin{aligned} &\inf{\cal S}(x),&&\text{if }x<z^{\gamma};\\ &\frac{\bar{s}_{i}+\gamma\bar{x}_{i}}{2-\gamma}\wedge\bar{s}_{i},&&\text{if }x=z^{\gamma};\\ &\sup{\cal S}(x),&&\text{if }x>z^{\gamma}.\end{aligned}\right$. (34)
If $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$, Assumption 1(ii) and (33) lead to the following relationship:
$\frac{\partial}{\partial s}g_{i}^{\gamma}\big(z^{\gamma},s(z^{\gamma})\big)$ $=\frac{\partial}{\partial s}g_{i}^{\gamma}(\tilde{x}_{i}^{\gamma},\bar{s}_{i})$
$=\tilde{x}_{i}^{\gamma}Q^{\prime}\left(\tilde{x}_{i}^{\gamma}+\bar{s}_{i}\right)-\left(1-\gamma\right)\left(\bar{x}_{i}-\tilde{x}_{i}^{\gamma}\right)Q^{\prime}\left(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(\tilde{x}_{i}^{\gamma}+\bar{s}_{i})\right)$
$=Q\left(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)\left(\tilde{x}_{i}^{\gamma}+\bar{s}_{i}\right)\right)-Q\left(\tilde{x}_{i}^{\gamma}+\bar{s}_{i}\right)$
$\leq 0,$
where the third equality stems from (11) and the inequality holds by observing that $\tilde{x}_{i}^{\gamma}\leq\bar{x}_{i}/2-(1-\gamma)\bar{s}_{i}/[2(2-\gamma)]<({\bar{x}_{i}-(1-\gamma)\bar{s}_{i}})/({2-\gamma})$ if $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$. On the other hand, if $\gamma\leq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$, then (33) suggests that
$\frac{\partial}{\partial s}g_{i}^{\gamma}\big(z^{\gamma},s(z^{\gamma})\big)$ $=\frac{\partial}{\partial s}g_{i}^{\gamma}\left(\frac{(1-\gamma)\bar{x}_{i}}{2-\gamma},\frac{\bar{s}_{i}+\gamma\bar{x}_{i}}{2-\gamma}\right)$ (35)
$=\frac{(1-\gamma)\bar{x}_{i}}{2-\gamma}Q^{\prime}\left(\frac{\bar{x}_{i}+\bar{s}_{i}}{2-\gamma}\right)-\frac{(1-\gamma)\bar{x}_{i}}{2-\gamma}Q^{\prime}\left(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)\frac{\bar{x}_{i}+\bar{s}_{i}}{2-\gamma}\right)$
$=0.$
These results imply that $s(z^{\gamma})\in{\cal S}(z^{\gamma})$ for all $\gamma\in[0,1]$ since $g_{i}^{\gamma}(x,\cdot)$ is convex for each $x\in[0,\bar{x}_{i}]$, and consequently, $s(x)\in{\cal S}(x)$ for all $x\in[0,\bar{x}_{i}]$ by the continuity of $g_{i}^{\gamma}(x,\cdot)$ on $0,\bar{s}_{i}]$ for each $x\in[0,\bar{x}_{i}]$.
Furthermore, by Assumption 1(ii) and (21d), the following holds for all $(x,s)\in[0,\bar{x}_{i}]\times[0,\bar{s}_{i}]$:
$\frac{\partial^{2}g_{i}^{\gamma}(x,s)}{\partial x\partial s}$ $=Q^{\prime}(x+s)+xQ^{\prime\prime}(x+s)+(1-\gamma)Q^{\prime}\left(\bar{x}_{i}-x+\bar{s}_{i}-s-\gamma(x+s)\right)$ (36)
$~~~~+(1-\gamma)^{2}(\bar{x}_{i}-x)Q^{\prime\prime}\left(\bar{x}_{i}-x+\bar{s}_{i}-s-\gamma(x+s)\right)$
$\leq Q^{\prime}(x+s)+xQ^{\prime\prime}(x+s)+(1-\gamma)Q^{\prime}\left(\bar{x}_{i}-x+\bar{s}_{i}-s-\gamma(x+s)\right)$
$~~~~+(1-\gamma)(\bar{x}_{i}-x)Q^{\prime\prime}\left(\bar{x}_{i}-x+\bar{s}_{i}-s-\gamma(x+s)\right)$
$<0.$
Then, by Lemma 1 of Amir (2005), $g_{i}^{\gamma}(x,s)$ is submodular in $0,\bar{x}_{i}]\times[0,\bar{s}_{i}]$, i.e.,
$g_{i}^{\gamma}(x,s)+g_{i}^{\gamma}(x^{\prime},s^{\prime})\geq g_{i}^{\gamma}(x\wedge x^{\prime},s\wedge s^{\prime})+g_{i}^{\gamma}(x\vee x^{\prime},s\vee s^{\prime})~~\text{for all}~(x^{\prime},s^{\prime})\in[0,\bar{x}_{i}]\times[0,\bar{s}_{i}].$
Thus, Topkis’s Monotonicity Theorem (see, e.g., Amir, 2005, Theorem 1) implies that $\inf{\cal S}(x)$ and $\sup{\cal S}(x)$ are non-decreasing in $x$. This leads to the following relationship:
$\begin{cases}s(x)=\inf{\cal S}(x)\leq\inf{\cal S}(z^{\gamma})\leq s(z^{\gamma}),~~&\text{if }x<z^{\gamma};\\ s(x)=\sup{\cal S}(x)\geq\sup{\cal S}(z^{\gamma})\geq s(z^{\gamma})~~&\text{if }x>z^{\gamma},\end{cases}$ (37)
which demonstrates that $s(x)$ is non-decreasing in $x$.
Step 3. We finally prove $x^{\gamma}_{i}=z^{\gamma}$ by showing that $g_{i}^{\gamma}(\cdot,s(\cdot))$ is strictly increasing on $0,z^{\gamma})$ and is strictly decreasing on $(z^{\gamma},\bar{x}_{i}]$, which immediately implies that $x^{\gamma}_{i}=z^{\gamma}\leq\bar{x}_{i}/2$ for any $\gamma\in[0,1]$ by (33) and the definition of $z^{\gamma}$. We observe that for any $x,\tilde{x}$ with $x<\tilde{x}$,
$g_{i}^{\gamma}\big(\tilde{x},s(\tilde{x})\big)-g_{i}^{\gamma}\big(x,s(x)\big)$ (38)
[…]
$~~~~-xQ\big(x+s(x)\big)-(\bar{x}_{i}-x)Q\Big(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)\big(x+s(x)\big)\Big)$
$))\displaystyle=(\tilde{x}-x)\bigg(Q\big(\tilde{x}+s(\tilde{x})\big)-Q\Big(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)\big(\tilde{x}+s(\tilde{x}$$\big)\Big$$\bigg)$
$~~~~+x\Big(Q\big(\tilde{x}+s(\tilde{x})\big)-Q\big(x+s(x)\big)\Big)$
$~~~~+(\bar{x}_{i}-x)\bigg(Q\Big(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)\big(\tilde{x}+s(\tilde{x})\big)\Big)-Q\Big(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)\big(x+s(x)\big)\Big)\bigg).$
We first fix $y,\tilde{y}\in[0,z^{\gamma})$ with $y<\tilde{y}$. We assume that $s(y)<\bar{s}_{i}$. Then, by (38), there exist $)\theta_{1}\in(y+s(y),\tilde{y}+s(\tilde{y}$)) and $\theta_{2}\in(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(\tilde{y}+s(\tilde{y})),\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(y+s(y)))$ such that
$g_{i}^{\gamma}(\tilde{y},s(\tilde{y}))-g_{i}^{\gamma}(y,s(y))$ (39)
$))\displaystyle=(\tilde{y}-y)\bigg(Q\big(\tilde{y}+s(\tilde{y})\big)-Q\Big(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)\big(\tilde{y}+s(\tilde{y}$$\big)\Big$$\bigg)$ (40)
$~~~~+\Delta_{0}\big(yQ^{\prime}(\theta_{1})-(1-\gamma)(\bar{x}_{i}-y)Q^{\prime}(\theta_{2})\big)$ (41)
$>\Delta_{0}\big(yQ^{\prime}(\theta_{1})-(1-\gamma)(\bar{x}_{i}-y)Q^{\prime}(\theta_{2})\big)$ (42)
$\geq\Delta_{0}\big(yQ^{\prime}(\theta_{1})-(1-\gamma)(\bar{x}_{i}-y)Q^{\prime}(\theta_{2})\big)-\Delta_{0}\frac{\partial}{\partial s}g_{i}^{\gamma}(y,s(y))$ (43)
$=\Delta_{0}y\Big(Q^{\prime}(\theta_{1})-Q^{\prime}\big(y+s(y)\big)\Big)$
$~~~~+\Delta_{0}(1-\gamma)(\bar{x}_{i}-y)\bigg(Q^{\prime}(u_{0})-Q^{\prime}(\theta_{2})\bigg)$
$\geq 0,$ (44)
where $\Delta_{0}\coloneqq\tilde{y}+s(\tilde{y})-(y+s(y))>0$ and $u_{0}\coloneqq\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(y+s(y))$, (41) follows from the mean value theorem, (42) is satisfied since $Q$ is strictly decreasing by Assumption 1(ii) and $\tilde{y}+s(\tilde{y})<z^{\gamma}+s(z^{\gamma})\leq(\bar{x}_{i}+\bar{s}_{i})/(2-\gamma)$ by (33), (43) holds because $g_{i}^{\gamma}(y,\cdot)$ is convex and $s(y)$ is its minimizer with $s(y)<\bar{s}_{i}$, and (44) stems from the convexity of $Q$ in Assumption 1(iv). On the other hand, if $\bar{s}_{i}, s(y)=s(y~)=$$s(\tilde{y})=\bar{s}_{i}\leq s(z^{\gamma})$ since $s(\cdot)$ is non-decreasing and $y<z^{\gamma}$. This implies that $\gamma\geq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$ and $(1-\gamma)\bar{x}_{i}/(2-\gamma)\leq\tilde{x}_{i}^{\gamma}$ by the definition of $s(\cdot)$ and (33), respectively. Hence, we have $y<\tilde{y}\leq z^{\gamma}=\tilde{x}_{i}^{\gamma}$. Furthermore, by (30) and the strict concavity of $g_{i}^{\gamma}(\cdot,\bar{s}_{i})$, we have
$\frac{\partial g_{i}^{\gamma}}{\partial x}(x,s(x))=\frac{\partial g_{i}^{\gamma}}{\partial x}(x,\bar{s}_{i})>\frac{\partial g_{i}^{\gamma}}{\partial x}(\tilde{x}_{i}^{\gamma},\bar{s}_{i})=0~~\text{for any }x\in[y,z^{\gamma}).$ (45)
This indicates that $)g_{i}^{\gamma}(y,s(y))=g_{i}^{\gamma}(y,\bar{s}_{i})<g_{i}^{\gamma}(\tilde{y},\bar{s}_{i})=g_{i}^{\gamma}(\tilde{y},s(\tilde{y}$)).
We next fix $t,\tilde{t}\in(z^{\gamma},\bar{x}_{i}]$ with $t<\tilde{t}$. If $s(t)<\bar{s}_{i}$, we have $s(z^{\gamma})={(\bar{s}_{i}+\gamma\bar{x}_{i})}/{(2-\gamma)}$ by (34) and $z^{\gamma}=(1-\gamma)\bar{x}_{i}/(2-\gamma)$ by (33). Then, by letting $\Delta_{1}\coloneqq\tilde{t}+s(\tilde{t})-(t+s(t))>0$, there exist $)\zeta_{1}\in(t+s(t),\tilde{t}+s(\tilde{t}$)) and $\zeta_{2}\in(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(\tilde{t}+s(\tilde{t})),\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(t+s(t)))$ such that
$g_{i}^{\gamma}(\tilde{t},s(\tilde{t}))-g_{i}^{\gamma}(t,s(t))$ (46)
$))\displaystyle=(\tilde{t}-t)\bigg(Q\big(\tilde{t}+s(\tilde{t})\big)-Q\Big(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)\big(\tilde{t}+s(\tilde{t}$$\big)\Big$$\bigg)$ (47)
$~~~~+\Delta_{1}\big(tQ^{\prime}(\zeta_{1})-(1-\gamma)(\bar{x}_{i}-t)Q^{\prime}(\zeta_{2})\big)$ (48)
$<\Delta_{1}\big(tQ^{\prime}(\zeta_{1})-(1-\gamma)(\bar{x}_{i}-t)Q^{\prime}(\zeta_{2})\big)$ (49)
$<\Delta_{1}\big((1-\gamma)(\bar{x}_{i}-t)Q^{\prime}(\zeta_{1})-tQ^{\prime}(\zeta_{2})\big)$ (50)
$\leq\Delta_{1}\big((1-\gamma)(\bar{x}_{i}-t)Q^{\prime}(\zeta_{1})-tQ^{\prime}(\zeta_{2})\big)+\Delta_{1}\frac{\partial}{\partial s}g_{i}^{\gamma}(t,s(t))$ (51)
$=\Delta_{1}t\Big(Q^{\prime}(t+s(t))-Q^{\prime}\big(\zeta_{2}\big)\Big)$
$~~~~+\Delta_{1}(1-\gamma)(\bar{x}_{i}-t)\bigg(Q^{\prime}(\zeta_{1})-Q^{\prime}(u_{1})\bigg)$
$\leq 0,$ (52)
where $u_{1}\coloneqq\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(t+s(t))$, (48) is due to (38) and the mean value theorem, (49) is satisfied since $\tilde{t}+s(\tilde{t})>z^{\gamma}+s(z^{\gamma})=(\bar{x}_{i}+\bar{s}_{i})/(2-\gamma)$, (50) follows by the fact that $t>z^{\gamma}=(1-\gamma)\bar{x}_{i}/(2-\gamma)$ and $Q^{\prime}(\cdot)<0$, (51) holds because $g_{i}^{\gamma}(t,\cdot)$ is convex and $s(t)$ is its minimizer with $s(t)<\bar{s}_{i}$, and (52) stems from Assumption 1(iv) and the inequalities $\zeta_{2}<\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(t+s(t))<t+s(t)<\zeta_{1}$. On the other hand, if $s(t)=\bar{s}_{i}$, then for any $x\in(t,\bar{x}_{i}]$, we have $x>z^{\gamma}$. By (31) and (21d), if $\gamma\leq\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$,
$\frac{\partial g_{i}^{\gamma}}{\partial x}\left(\frac{(1-\gamma)\bar{x}_{i}}{2-\gamma},\bar{s}_{i}\right)\leq 0,$
and thus,
$\frac{\partial g_{i}^{\gamma}}{\partial x}(x,s(x))=\frac{\partial g_{i}^{\gamma}}{\partial x}(x,\bar{s}_{i})<\frac{\partial g_{i}^{\gamma}}{\partial x}(z^{\gamma},\bar{s}_{i})\leq 0$
by (30), the definition of $z^{\gamma}$, and the strict concavity of $g_{i}^{\gamma}(\cdot,\bar{s}_{i})$. This suggests that $)g_{i}^{\gamma}(t,s(t))=g_{i}^{\gamma}(t,\bar{s}_{i})>g_{i}^{\gamma}(\tilde{t},\bar{s}_{i})=g_{i}^{\gamma}(\tilde{t},s(\tilde{t}$)). Consequently, the result follows. ∎
Proof.
Proof of Theorem 2. Fix $i\in[m]$ and define $g_{i}^{\gamma}(x,s)\coloneqq e_{i}+xQ(x+s)+(\bar{x}_{i}-x)Q(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(x+s))$ for each $x\in[0,\bar{x}_{i}]$ and $s\in[0,\bar{s}_{i}]$. Then, (9) can be rewritten as
$\max_{x\in[0,\,\bar{x}_{i}]}\,\min_{s\in\left[0,\,\bar{s}_{i}\right]}\;g_{i}^{\gamma}(x,\,s).$ (53)
According to the proof of Theorem 1, we observe that if $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$, then $x^{\gamma}_{i}=\tilde{x}_{i}^{\gamma}$ and $s(x^{\gamma}_{i})\coloneqq\mathop{\arg\min}_{s\in[0,\bar{s}_{i}]}g_{i}^{\gamma}(x^{\gamma}_{i},s)=\bar{s}_{i}$. Furthermore, since $(\partial/\partial x)g_{i}^{\gamma}(\cdot,\bar{s}_{i})$ is strictly decreasing on $0,\bar{x}_{i}]$ by Assumption 1(iv) and (21c), it suffices to show that $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$ and $\bar{x}_{i}\leq\bar{s}_{i}$ imply
$\frac{\partial^{2}g_{i}}{\partial x\partial\gamma}(\tilde{x}_{i}^{\gamma},\bar{s}_{i})\geq 0.$ (54)
Assume that $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$ and $\bar{x}_{i}\leq\bar{s}_{i}$. Then, we have
$\frac{(1-\gamma^{2})\bar{x}_{i}}{2-\gamma}+\gamma\bar{s}_{i}\geq\bar{x}_{i}$
since it holds when $\gamma=\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$ and the left-hand side increases in $\gamma$. Thus, by (33), we obtain $(1+\gamma)\tilde{x}_{i}^{\gamma}+\gamma\bar{s}_{i}\geq\bar{x}_{i}$. This suggests that $\gamma\left(\tilde{x}_{i}^{\gamma}+\bar{s}_{i}\right)^{2}\geq\left(\tilde{x}_{i}^{\gamma}+\bar{s}_{i}\right)\left(\bar{x}_{i}-\tilde{x}_{i}^{\gamma}\right)\geq\left(\bar{x}_{i}-\tilde{x}_{i}^{\gamma}\right)^{2}$, and therefore, we observe that
$\left(2\tilde{x}_{i}^{\gamma}+\bar{s}_{i}-\bar{x}_{i}\right)\left(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\right)-(1-\gamma)\left(\tilde{x}_{i}^{\gamma}+\bar{s}_{i}\right)\left(\bar{x}_{i}-\tilde{x}_{i}^{\gamma}\right)$ (55)
$=\gamma\left(\tilde{x}_{i}^{\gamma}+\bar{s}_{i}\right)^{2}-\left(\bar{x}_{i}-\tilde{x}_{i}^{\gamma}\right)^{2}$
$\geq 0,$
which also gives $2\tilde{x}_{i}^{\gamma}+\bar{s}_{i}-\bar{x}_{i}\geq 0$. Consequently, we obtain the following relationship:
$\frac{\partial^{2}g_{i}}{\partial x\partial\gamma}(\tilde{x}_{i}^{\gamma},\bar{s}_{i})$ (56)
$=(\bar{x}_{i}-\bar{s}_{i}-2\tilde{x}_{i}^{\gamma})Q^{\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)-(1-\gamma)(\bar{x}_{i}-\tilde{x}_{i}^{\gamma})(\tilde{x}_{i}^{\gamma}+\bar{s}_{i})Q^{\prime\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)$
$\geq-(2\tilde{x}_{i}^{\gamma}+\bar{s}_{i}-\bar{x}_{i})\Big(Q^{\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)+\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)Q^{\prime\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)\Big)$
$\geq 0,$
where the first and second inequalities hold by (55) and Assumption 1(v), respectively. ∎
Proof.
Proof of Proposition 2. Fix $i\in[m]$ and $\bar{x}_{i}>0$. By Theorem 1, it suffices to prove that $x^{\gamma}_{i}$ decreases in $\gamma$ if $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$ and $\bar{s}_{i}$ is sufficiently small. Define $g_{i}^{\gamma}(x,s)\coloneqq e_{i}+xQ(x+s)+(\bar{x}_{i}-x)Q(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(x+s))$ for each $x\in[0,\bar{x}_{i}]$ and $s\in[0,\bar{s}_{i}]$. Fix $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$. We observe that if $\gamma>\sqrt{3}-1$, then for some $\theta\in[(\bar{x}_{i}+\bar{s}_{i})/2,(1+\gamma)(\bar{x}_{i}+\bar{s}_{i})/2]$,
$\frac{\partial g_{i}}{\partial x}\bigg(\frac{\bar{x}_{i}-\bar{s}_{i}}{2},\bar{s}_{i}\bigg)$ (57)
$=Q\bigg(\frac{\bar{x}_{i}+\bar{s}_{i}}{2}\bigg)+\frac{\bar{x}_{i}-\bar{s}_{i}}{2}Q^{\prime}\bigg(\frac{\bar{x}_{i}+\bar{s}_{i}}{2}\bigg)-Q\bigg(\frac{(1+\gamma)(\bar{x}_{i}+\bar{s}_{i})}{2}\bigg)-\frac{(1-\gamma)(\bar{x}_{i}+\bar{s}_{i})}{2}Q^{\prime}\bigg(\frac{(1+\gamma)(\bar{x}_{i}+\bar{s}_{i})}{2}\bigg)$
$=\frac{\bar{x}_{i}-\bar{s}_{i}}{2}Q^{\prime}\bigg(\frac{\bar{x}_{i}+\bar{s}_{i}}{2}\bigg)-\frac{\gamma(\bar{x}_{i}+\bar{s}_{i})}{2}Q^{\prime}\bigg(\frac{\bar{x}_{i}+\bar{s}_{i}}{2}\bigg)-\frac{\gamma^{2}(\bar{x}_{i}+\bar{s}_{i})^{2}}{8}Q^{\prime\prime}(\theta)-\frac{(1-\gamma)(\bar{x}_{i}+\bar{s}_{i})}{2}Q^{\prime}\left(\frac{(1+\gamma)(\bar{x}_{i}+\bar{s}_{i})}{2}\right)$
$\leq\frac{\bar{x}_{i}-\bar{s}_{i}}{2}Q^{\prime}\bigg(\frac{\bar{x}_{i}+\bar{s}_{i}}{2}\bigg)-\frac{\gamma(\bar{x}_{i}+\bar{s}_{i})}{2}Q^{\prime}\bigg(\frac{\bar{x}_{i}+\bar{s}_{i}}{2}\bigg)-\frac{\gamma^{2}(\bar{x}_{i}+\bar{s}_{i})^{2}}{8}Q^{\prime\prime}(\theta)-\frac{(1-\gamma)(\bar{x}_{i}+\bar{s}_{i})}{2}Q^{\prime}\bigg(\frac{\bar{x}_{i}+\bar{s}_{i}}{2}\bigg)$
$=-\bar{s}_{i}Q^{\prime}\bigg(\frac{\bar{x}_{i}+\bar{s}_{i}}{2}\bigg)-\frac{\gamma^{2}(\bar{x}_{i}+\bar{s}_{i})^{2}}{8}Q^{\prime\prime}(\theta)$
$\leq-Q^{\prime}\bigg(\frac{\bar{x}_{i}+\bar{s}_{i}}{2}\bigg)\bigg(\bar{s}_{i}+\frac{(\sqrt{3}-1)^{2}\bar{x}_{i}^{2}\delta}{8Q^{\prime}(0)}\bigg)$
$=-Q^{\prime}\bigg(\frac{\bar{x}_{i}+\bar{s}_{i}}{2}\bigg)\bigg(\bar{s}_{i}+\frac{(2-\sqrt{3})\bar{x}_{i}^{2}\delta}{4Q^{\prime}(0)}\bigg),$
where $\delta=\min_{x\in[0,\bar{x}_{i}+\bar{s}_{i}]}Q^{\prime\prime}(x)>0$, the second equality follows from the first-order Taylor expansion, and the two inequalities hold since $Q^{\prime}(0)\leq Q^{\prime}((\bar{x}_{i}+\bar{s}_{i})/2)\leq Q^{\prime}((1+\gamma)(\bar{x}_{i}+\bar{s}_{i})/2)\leq 0$ by Assumption 1.
As in the proof of Theorem 1, define $\tilde{x}_{i}^{\gamma}$ as the solution of (11) with respect to $x$ given $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$. In the remainder of the proof, we assume that $\bar{s}_{i}<-{(2-\sqrt{3})\bar{x}_{i}^{2}\delta}/({4Q^{\prime}(0)})$ and claim that
$\frac{\partial^{2}g_{i}}{\partial x\partial\gamma}(\tilde{x}_{i}^{\gamma},\bar{s}_{i})=(\bar{x}_{i}-\bar{s}_{i}-2\tilde{x}_{i}^{\gamma})Q^{\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)-(1-\gamma)(\bar{x}_{i}-\tilde{x}_{i}^{\gamma})(\tilde{x}_{i}^{\gamma}+\bar{s}_{i})Q^{\prime\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)\leq 0.$ (58)
This is trivial if $\tilde{x}_{i}^{\gamma}<(\bar{x}_{i}-\bar{s}_{i})/2$. Thus, suppose that $\tilde{x}_{i}^{\gamma}\geq(\bar{x}_{i}-\bar{s}_{i})/2$. Then, since $(\partial/\partial x)g_{i}^{\gamma}(\cdot,\bar{s}_{i})$ is decreasing on $0,\bar{x}_{i}]$ and $(\partial/\partial x)g_{i}^{\gamma}(\tilde{x}_{i}^{\gamma},\bar{s}_{i})=0$ by the proof of Theorem 1, (33) and (57) suggests $\gamma\leq\sqrt{3}-1$ and
$\frac{\bar{x}_{i}-\bar{s}_{i}}{2}\leq\tilde{x}_{i}^{\gamma}\leq\frac{\bar{x}_{i}}{2}-\frac{(1-\gamma)\bar{s}_{i}}{2(2-\gamma)},$ (59)
We therefore obtain the following result:
$\frac{\partial^{2}g_{i}}{\partial x\partial\gamma}(\tilde{x}_{i}^{\gamma},\bar{s}_{i})$ $\leq-\frac{\bar{s}_{i}}{2-\gamma}Q^{\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)-(1-\gamma)\bigg(\frac{\bar{x}_{i}}{2}+\frac{(1-\gamma)\bar{s}_{i}}{2(2-\gamma)}\bigg)\frac{(\bar{x}_{i}+\bar{s}_{i})}{2}Q^{\prime\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)$ (60)
$\leq-\bar{s}_{i}Q^{\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)-(2-\sqrt{3})\frac{\bar{x}_{i}^{2}}{4}Q^{\prime\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)$
$=-Q^{\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)\bigg(\bar{s}_{i}+(2-\sqrt{3})\frac{\bar{x}_{i}^{2}}{4}\frac{Q^{\prime\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)}{Q^{\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)}\bigg)$
$\leq-Q^{\prime}\big(\bar{x}_{i}+\gamma\bar{s}_{i}-(1-\gamma)\tilde{x}_{i}^{\gamma}\big)\bigg(\bar{s}_{i}+\frac{(2-\sqrt{3})\bar{x}_{i}^{2}\delta}{4Q^{\prime}(0)}\bigg)$
$<0,$
where the first two inequalities comes from (59) and the third inequality holds by Assumption 1. Consequently, the above claim is demonstrated. This leads to the desired result because the proof of Theorem 1 shows that $(\partial/\partial x)g_{i}^{\gamma}(\cdot,\bar{s}_{i})$ is strictly decreasing on $0,\bar{x}_{i}]$ and that $x^{\gamma}_{i}=\tilde{x}_{i}^{\gamma}$ and $s(x^{\gamma}_{i})\coloneqq\mathop{\arg\min}_{s\in[0,\bar{s}_{i}]}g_{i}^{\gamma}(x^{\gamma}_{i},s)=\bar{s}_{i}$. ∎
Appendix B Proofs of Theoretical Results in Section 5
Throughout this section, we denote the 1-norm of a vector ${\bf u}\in\mathbb{R}^{d}$ by $\|{\bf u}\|_{1}\coloneqq\sum_{k=1}^{d}|u_{k}|$, and we define $\bar{p}^{\bf A}\coloneqq Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})$.
Proof.
Proof of Lemma 2. To facilitate the proof, we use the following lemma, whose proof is provided in Appendix C.
Lemma 3.
$\bar{p}^{\bf A}\leq p_{2}^{\bf A}$.
Since $\boldsymbol{\pi}(\cdot)$ is increasing and $p_{1}\coloneqq Q({\bf 1}^{\top}{\bf x})\geq Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})=\bar{p}^{\bf A}$, Lemma 3 implies that
$\bar{\boldsymbol{\pi}}^{\bf A}=\boldsymbol{\pi}({\bf z}\bar{p}^{\bf A})\leq\boldsymbol{\pi}({\bf x}p_{1}+({\bf z}-{\bf x})p_{2}^{\bf A})=\boldsymbol{\pi}^{\bf A}.$
Consequently, the following holds for all $j$:
$\psi_{j}^{\bf A}({\bf x})$ $=\frac{\left(L_{j}-e_{j}-x_{j}p_{1}-{\bf A}_{j}^{\top}\boldsymbol{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)$ (61)
$\leq\frac{\left(L_{j}-e_{j}-x_{j}p_{1}-{\bf A}_{j}^{\top}\bar{\boldsymbol{\pi}}^{\bf A}\right)^{+}}{\bar{p}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)$
$\leq\frac{\left(L_{j}-e_{j}-x_{j}\bar{p}^{\bf A}-{\bf A}_{j}^{\top}\bar{\boldsymbol{\pi}}^{\bf A}\right)^{+}}{\bar{p}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)$
$=\left(\frac{L_{j}-e_{j}-{\bf A}_{j}^{\top}\bar{\boldsymbol{\pi}}^{\bf A}}{\bar{p}^{\bf A}}-x_{j}\right)\wedge\left(z_{j}-x_{j}\right)$
$=\bar{x}^{\bf A}_{j}-x_{j},$
where the second last equality holds since $x_{j}\leq\bar{x}^{\bf A}_{j}\leq(L_{j}-e_{j}-{\bf A}_{j}^{\top}\bar{\boldsymbol{\pi}}^{\bf A})/\bar{p}^{\bf A}$ for all $j$. This completes the proof. ∎
Proof.
Proof of Proposition 3. To explicitly show the dependence of $\boldsymbol{\pi}^{\bf A}$ on the first-period liability repayment $\bf x$}, we write $\boldsymbol{\pi}^{\bf A}=\boldsymbol{\pi}^{\bf A}_{\bf x}$ in this proof. We first show that
$\boldsymbol{\pi}^{\bf A}_{\bf x}\geq\boldsymbol{\pi}({\bf c}({\bf x}))~~\text{for all}~{\bf x}\leq\bar{{\bf x}}^{\bf A}.$ (62)
Define ${\bf\psi}^{\bf A}({\bf x})\coloneqq(\psi_{1}^{\bf A}({\bf x}),\dots,\psi_{m}^{\bf A}({\bf x}))^{\top}$. Let $\boldsymbol{\phi}_{0}={\bf L}\wedge\left({\bf e}+{\bf x}Q({\bf 1}^{\top}{\bf x})+(\bar{{\bf x}}^{\bf A}-{\bf x})Q\big({\bf 1}^{\top}(\bar{{\bf x}}^{\bf A}-{\bf x})\big)+{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf x}\right)$ and $\boldsymbol{\phi}_{k+1}={\bf L}\wedge\left({\bf e}+{\bf x}Q({\bf 1}^{\top}{\bf x})+(\bar{{\bf x}}^{\bf A}-{\bf x})Q\big({\bf 1}^{\top}(\bar{{\bf x}}^{\bf A}-{\bf x})\big)+{\bf A}^{\top}\boldsymbol{\phi}_{k}\right)\text{ for }k=0,1,2,\dots.$ Then, $\boldsymbol{\pi}^{\bf A}_{\bf x}\geq\boldsymbol{\phi}_{0}$ since for any $j\in[m]$,
$\pi_{{\bf x},j}^{\bf A}$ $=L_{j}\wedge\left(e_{j}+x_{j}Q({\bf 1}^{\top}{\bf x})+(z_{j}-x_{j})Q\big({\bf 1}^{\top}{\bf\psi}^{\bf A}({\bf x})\big)+{\bf A}_{j}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf x}\right)$
$\geq L_{j}\wedge\left(e_{j}+x_{j}Q({\bf 1}^{\top}{\bf x})+(\bar{x}^{\bf A}_{j}-x_{j})Q\big({\bf 1}^{\top}(\bar{{\bf x}}^{\bf A}-{\bf x})\big)+{\bf A}_{j}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf x}\right)$
$=\phi_{0,j},$
where the inequality follows from Lemma 2 and the fact that $\bar{{\bf x}}^{\bf A}\leq{\bf z}$. Therefore, $\boldsymbol{\pi}^{\bf A}_{\bf x}\geq\boldsymbol{\phi}_{0}$. Furthermore, using a similar argument, it can be shown that if $\boldsymbol{\pi}^{\bf A}_{\bf x}\geq\boldsymbol{\phi}_{k}$ for some $k\geq 0$, then $\boldsymbol{\pi}^{\bf A}_{\bf x}\geq\boldsymbol{\phi}_{k+1}$. Hence, $\boldsymbol{\pi}_{\bf x}^{\bf A}\geq\boldsymbol{\phi}_{k}$ for all $k$ by induction, which suggests (62) since $\boldsymbol{\phi}_{k}\rightarrow\boldsymbol{\pi}({\bf c}({\bf x}))$ as $k$ increases.
Next, we prove that $\phi^{\bf A}_{i}(x_{i},\bar{{\bf x}}^{\bf A}_{-i})\leq L_{i}$ for any $x_{i}\leq\bar{x}^{\bf A}_{i}$. Define a vector ${\bf u}\in\mathbb{R}^{m}$ with $u_{i}=x_{i}$ and ${\bf u}_{-i}=\bar{{\bf x}}^{\bf A}_{-i}$. Note that $\psi_{j}^{\bf A}({\bf u})=0$ for all $j\neq i$ by Lemma 2. If $\psi_{i}^{\bf A}({\bf u})>0$, then $L_{i}-e_{i}-x_{i}Q({\bf 1}^{\top}{\bf u})-{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf u}>0$ by (16), and hence,
$\phi^{\bf A}_{i}(x_{i},\bar{{\bf x}}^{\bf A}_{-i})$ $=e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+\psi_{i}^{\bf A}({\bf u})Q(\psi_{i}^{\bf A}({\bf u}))+{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf u}$ (63)
$=\Big(L_{i}\vee\big(e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf u}\big)\Big)\wedge\Big(e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+(z_{i}-x_{i})Q(\psi_{i}^{\bf A}({\bf u}))+{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf u}\Big)$
$=L_{i}\wedge\Big(e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+(z_{i}-x_{i})Q(\psi_{i}^{\bf A}({\bf u}))+{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf u}\Big)$
$\leq L_{i},$
where the second equality holds by (16). If $\psi_{i}^{\bf A}({\bf u})=0$, then $\boldsymbol{\pi}^{\bf A}_{\bf u}=\boldsymbol{\pi}({\bf u}Q({\bf 1}^{\top}{\bf u}))$ by (15) and (16), and therefore,
$\phi^{\bf A}_{i}(x_{i},\bar{{\bf x}}^{\bf A}_{-i})$ $=e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+{\bf A}^{\top}\boldsymbol{\pi}({\bf u}Q({\bf 1}^{\top}{\bf u}))$
$\leq e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf u}Q({\bf 1}^{\top}{\bf u}))-\left(e_{i}+\bar{x}^{\bf A}_{i}\bar{p}^{\bf A}+{\bf A}_{i}^{\top}\bar{\boldsymbol{\pi}}^{\bf A}\right)+L_{i}$
$=x_{i}Q({\bf 1}^{\top}{\bf u})-\bar{x}^{\bf A}_{i}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})+{\bf A}_{i}^{\top}\big(\boldsymbol{\pi}({\bf u}Q({\bf 1}^{\top}{\bf u}))-\boldsymbol{\pi}(\bar{{\bf x}}^{\bf A}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A}))\big)+L_{i}$
$\leq x_{i}Q({\bf 1}^{\top}{\bf u})-\bar{x}^{\bf A}_{i}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})+{\bf A}_{i}^{\top}\Big(\boldsymbol{\pi}\big(({\bf u}Q({\bf 1}^{\top}{\bf u}))\vee(\bar{{\bf x}}^{\bf A}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A}))\big)-\boldsymbol{\pi}(\bar{{\bf x}}^{\bf A}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A}))\Big)+L_{i}$
$\leq x_{i}Q({\bf 1}^{\top}{\bf u})-\bar{x}^{\bf A}_{i}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})+\big\|({\bf u}Q({\bf 1}^{\top}{\bf u}))\vee(\bar{{\bf x}}^{\bf A}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A}))-\bar{{\bf x}}^{\bf A}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})\big\|_{1}$
$=x_{i}Q({\bf 1}^{\top}{\bf u})-\bar{x}^{\bf A}_{i}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})+({\bf 1}^{\top}\bar{{\bf x}}^{\bf A}_{-i})\big(Q({\bf 1}^{\top}{\bf u})-Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})\big)+L_{i}$
$=({\bf 1}^{\top}{\bf u})Q({\bf 1}^{\top}{\bf u})-({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})+L_{i}$
$\leq L_{i},$
where the first inequality comes from the definition of $\bar{x}^{\bf A}_{i}$ and the assumption that $\bar{{\bf x}}^{\bf A}>{\bf 0}$, the second inequality holds because $\boldsymbol{\pi}(\cdot)$ is increasing, the third inequality follows since $\boldsymbol{\pi}(\cdot)$ is non-expansive and ${\bf A}_{i}<{\bf 1}$, the third equality is satisfied as ${\bf u}_{-i}Q({\bf 1}^{\top}{\bf u})=\bar{{\bf x}}^{\bf A}_{-i}Q(x_{i}+{\bf 1}^{\top}\bar{{\bf x}}^{\bf A}_{-i})\geq\bar{{\bf x}}^{\bf A}_{-i}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})$ by Assumption 1(ii) and $u_{i}Q({\bf 1}^{\top}{\bf u})=x_{i}Q(x_{i}+{\bf 1}^{\top}\bar{{\bf x}}^{\bf A}_{-i})\leq\bar{x}^{\bf A}_{i}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})$ by (21a), and the last inequality stems from Assumption 1(iii).
Finally, we demonstrate the two statements. Define ${\bf v}\in\mathbb{R}^{m}$ where $v_{i}=x_{i}$ and ${\bf v}_{-i}=\mathop{\arg\min}_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\phi^{\bf A}_{i}(x_{i},{\bf x}_{-i})$. Note that $\phi^{\bf A}_{i}(x_{i},{\bf v}_{-i})\leq\phi^{\bf A}_{i}(x_{i},\bar{{\bf x}}^{\bf A}_{-i})\leq L_{i}$. Assume that $\phi^{\bf A}_{i}(x_{i},{\bf v}_{-i})<L_{i}$. Then, by (17), we get
$\psi_{i}^{\bf A}({\bf v})=\frac{\phi^{\bf A}_{i}(x_{i},{\bf v}_{-i}^{\bf A})-e_{i}-x_{i}Q\big({\bf 1}^{\top}{\bf v}\big)-{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf v}}{Q\left(\sum_{j=1}^{m}\psi_{j}^{\bf A}({\bf v})\right)}<\frac{\left(L_{i}-e_{i}-x_{i}Q\big({\bf 1}^{\top}{\bf v}\big)-{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf v}\right)^{+}}{Q\left(\sum_{j=1}^{m}\psi_{j}^{\bf A}({\bf v})\right)},$ (64)
which implies that $\psi_{i}^{\bf A}({\bf v})=z_{i}-x_{i}$ by (16). Therefore, we obtain
$\phi^{\bf A}_{i}(x_{i},{\bf v}_{-i})$ $=e_{i}+x_{i}Q\big({\bf 1}^{\top}{\bf v}\big)+(z_{i}-x_{i})Q\Big(\sum_{j=1}^{m}\psi_{j}^{\bf A}({\bf v})\Big)+{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf v}$ (65)
$\geq e_{i}+x_{i}Q\big({\bf 1}^{\top}{\bf v}\big)+(\bar{x}^{\bf A}_{i}-x_{i})Q\Big(\sum_{j=1}^{m}(\bar{x}^{\bf A}_{j}-v_{j})\Big)+{\bf A}_{i}^{\top}\boldsymbol{\pi}\big({\bf c}({\bf v})\big)$
$=\Phi^{\bf A}_{i}(x_{i},{\bf v}_{-i}),$
where the inequality holds by (62), Lemma 2, and the fact that $\bar{x}^{\bf A}_{i}\leq z_{i}$. Consequently, $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\Phi^{\bf A}_{i}(x_{i},{\bf x}_{-i})<L_{i}$, which demonstrates the first statement by contrapositive.
The second statement is straightforward if $\phi^{\bf A}_{i}(x_{i},{\bf v}_{-i}^{\bf A})=L_{i}$. Otherwise, it holds by (65). ∎
Proof.
Proof of Theorem 3. Fix $i\in[m]$. If $\bar{x}^{\bf A}_{i}=0$, the statement is straightforward. We assume that $\bar{x}^{\bf A}_{i}>0$. Note that $\Phi^{\bf A}_{i}(x,{\bf x}_{-i})=\Phi^{\bf A}_{i}(\bar{x}^{\bf A}_{i}-x,\bar{{\bf x}}^{\bf A}_{-i}-{\bf x}_{-i})$ for all $x\in[0,\bar{x}^{\bf A}_{i}]$. Thus, it suffices to show that for all $x\in[0,\bar{x}^{\bf A}_{i}/2)$,
$\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}\left(\frac{\bar{x}^{\bf A}_{i}}{2},{\bf x}_{-i}\right)>\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}\left(x,{\bf x}_{-i}\right)$ (66)
We prove (66) in the following four steps.
Step 1. Let $\bar{\zeta}=\sum_{j=1}^{m}\bar{x}^{\bf A}_{j}$ and $\bar{s}^{\bf A}_{i,j}=\bar{\zeta}-\bar{x}^{\bf A}_{i}-\bar{x}^{\bf A}_{j}$ for $j\in[m]$. For all $y\in[0,\bar{x}^{\bf A}_{i}/2]$, $j\neq i$, and $\zeta\in{\cal Z}(y)\coloneqq[y,\bar{\zeta}-\bar{x}^{\bf A}_{i}+y]$, define
$\mu_{j}\left(y,\zeta\right)$ $\coloneqq\left(\zeta-y-\bar{s}^{\bf A}_{i,j}\right)^{+}\left(Q(\zeta)-Q(\bar{\zeta}-\zeta)\right)+\bar{x}^{\bf A}_{j}Q(\bar{\zeta}-\zeta),$ (67)
$\nu_{j}\left(y,\zeta\right)$ $\coloneqq\left(\bar{x}^{\bf A}_{j}\wedge\left(\zeta-y\right)\right)\left(Q(\zeta)-Q(\bar{\zeta}-\zeta)\right)+\bar{x}^{\bf A}_{j}Q(\bar{\zeta}-\zeta),$ (68)
and
${\cal B}(y,\zeta)\coloneqq\left\{{\bf c}\in\mathbb{R}^{m}\Bigg|\begin{array}[]{l}{\bf 1}^{\top}{\bf c}=\zeta\left(Q(\zeta)-Q(\bar{\zeta}-\zeta)\right)+\bar{\zeta}Q(\bar{\zeta}-\zeta),c_{i}=y\left(Q(\zeta)-Q(\bar{\zeta}-\zeta)\right)+\bar{x}^{\bf A}_{i}Q(\bar{\zeta}-\zeta),\\ c_{j}\in\left[\mu_{j}(y,\zeta)\wedge\nu_{j}(y,\zeta),\mu_{j}(y,\zeta)\vee\nu_{j}(y,\zeta)\right],~\forall j\neq i\end{array}\right\}.$
For each $y\in[0,\bar{x}^{\bf A}_{i}/2]$, $j\neq i$, and $\zeta\in{\cal Z}(y)$, $(\zeta-y-\bar{s}^{\bf A}_{i,j})^{+}\leq\bar{x}^{\bf A}_{j}\wedge(\zeta-y)$, and thus, we have
$\begin{cases}\mu_{j}(y,\zeta)\leq\nu_{j}(y,\zeta),&\text{if }\zeta\leq\bar{\zeta}/2;\\ \mu_{j}(y,\zeta)>\nu_{j}(y,\zeta),&\text{otherwise}.\end{cases}$ (69)
Observe that
$\sum_{j\neq i}\left(\zeta-y-\bar{s}^{\bf A}_{i,j}\right)^{+}\leq\zeta-y\leq\sum_{j\neq i}\left(\bar{x}^{\bf A}_{j}\wedge\left(\zeta-y\right)\right),$ (70)
where the first inequality holds because if $\zeta-y>\bar{s}^{\bf A}_{i,k^{*}}$, where $k^{*}\coloneqq\mathop{\arg\min}_{k\neq i}\bar{s}^{\bf A}_{i,k}$,
$\sum_{j\neq i}(\zeta-y-\bar{s}^{\bf A}_{i,j})^{+}\leq\zeta-y-\bar{s}^{\bf A}_{i,k^{*}}+\sum_{j\neq i,k^{*}}\bar{x}^{\bf A}_{j}=\zeta-y,$
and $\sum_{j\neq i}(\zeta-y-\bar{s}^{\bf A}_{i,j})^{+}=0\leq\zeta-y$ otherwise. Accordingly, if $\zeta>\bar{\zeta}/2$, we obtain
$\sum_{j\neq i}\mu_{j}(y,\zeta)\leq(\zeta-y)\left(Q(\zeta)-Q(\bar{\zeta}-\zeta)\right)+(\bar{\zeta}-\bar{x}^{\bf A}_{i})Q(\bar{\zeta}-\zeta)\leq\sum_{j\neq i}\nu_{j}(y,\zeta),$
and otherwise
$\sum_{j\neq i}\mu_{j}(y,\zeta)\geq(\zeta-y)\left(Q(\zeta)-Q(\bar{\zeta}-\zeta)\right)+(\bar{\zeta}-\bar{x}^{\bf A}_{i})Q(\bar{\zeta}-\zeta)\geq\sum_{j\neq i}\nu_{j}(y,\zeta).$
This implies that ${\cal B}(y,\zeta)$ is non-empty for any $y\in[0,\bar{x}^{\bf A}_{i}/2]$ and $\zeta\in{\cal Z}(y)$.
Step 2. For $y\in[0,\bar{x}^{\bf A}_{i}/2]$, we claim that
$\tilde{g}_{i}(y)\coloneqq\min_{\zeta\in{\cal Z}(y)}\min_{{\bf c}\in{\cal B}(y,\zeta)}\left\{e_{i}+c_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf c})\right\}=\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(y,{\bf x}_{-i}).$ (71)
We first prove that $\tilde{g}_{i}(y)\leq\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(y,{\bf x}_{-i})$. For any ${\bf x}^{0}$ with $x_{i}^{0}=y$ and ${\bf x}_{-i}^{0}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]$, let $\zeta^{0}={\bf 1}^{\top}{\bf x}^{0}$ and ${\bf c}^{0}={\bf c}({\bf x}^{0})$. Then, by (18), for all $j\in[m]$,
$c_{j}^{0}=x_{j}^{0}\left(Q(\zeta^{0})-Q(\bar{\zeta}-\zeta^{0})\right)+\bar{x}^{\bf A}_{j}Q(\bar{\zeta}-\zeta^{0}).$ (72)
Obviously, $\zeta^{0}\in{\cal Z}(y)$. Moreover, $(\zeta^{0}-\bar{s}^{\bf A}_{i,j}-y)^{+}\leq x_{j}^{0}\leq\zeta^{0}-y$, which implies that for all $j\neq i$,
$\mu_{j}(y,\zeta)\wedge\nu_{j}(y,\zeta)\leq c_{j}^{0}\leq\mu_{j}(y,\zeta)\vee\nu_{j}(y,\zeta).$
by (67), (68) and (72). Therefore, we have ${\bf c}^{0}\in{\cal B}(y,\zeta^{0})$. Consequently, since $\Phi^{\bf A}_{i}(y,{\bf x}_{-i}^{0})=e_{i}+c_{i}^{0}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf c}^{0})$ and $\tilde{g}_{i}(y)\leq e_{i}+c_{i}^{0}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf c}^{0})$, $\tilde{g}_{i}(y)\leq\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(y,{\bf x}_{-i})$ holds.
We now show that $\tilde{g}_{i}(y)\geq\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(y,{\bf x}_{-i})$. Fix $\zeta^{1}\in{\cal Z}(y)$ and ${\bf c}^{1}\in{\cal B}(y,\zeta^{1})$. Assume that $\zeta^{1}\neq\bar{\zeta}/2$. Let ${\bf x}^{1}\in\mathbb{R}^{m}$ with $x_{i}^{1}=y$ and
$x_{j}^{1}=\frac{c_{j}^{1}-\bar{x}^{\bf A}_{j}Q(\bar{\zeta}-\zeta^{1})}{Q(\zeta^{1})-Q(\bar{\zeta}-\zeta^{1})},\quad j\neq i.$
It is easy to see that $x_{j}^{1}\in[(\zeta^{1}-\bar{s}^{\bf A}_{i,j}-y)^{+},\bar{x}^{\bf A}_{j}\wedge(\zeta^{1}-y)]\subset[0,\bar{x}^{\bf A}_{j}]$ for $j\neq i$, which implies that ${\bf x}_{-i}^{1}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]$. Further, it can be verified that ${\bf c}^{1}={\bf c}({\bf x}^{1})$ by (18). Otherwise, if $\zeta^{1}=\bar{\zeta}/2$, $\bar{\zeta}/2-y\leq\bar{\zeta}-\bar{x}^{\bf A}_{i}$ since $\zeta^{1}\in{\cal Z}(y)$. Then there exists some ${\bf x}^{1}\in[{\bf 0},\bar{{\bf x}}^{\bf A}]$ such that ${\bf 1}^{\top}{\bf x}_{-i}^{1}=\bar{\zeta}/2-y$, $x_{i}^{1}=y$, and $c_{j}({\bf x}^{1})=\bar{x}^{\bf A}_{j}Q(\bar{\zeta}/2)=c_{j}^{1}$ for all $j\in[m]$, implying that ${\bf c}^{1}={\bf c}({\bf x}^{1})$.
Hence, $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(y,{\bf x}_{-i})\leq e_{i}+c_{i}^{1}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf c}^{1})$ for any feasible $\zeta^{1}$ and ${\bf c}^{1}$, which implies that $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(y,{\bf x}_{-i})\leq\tilde{g}_{i}(y)$. As a consequence, (71) follows.
Step 3. We first fix
$\tilde{\zeta}\in\mathop{\arg\min}_{\zeta\in{\cal Z}\left(\bar{x}^{\bf A}_{i}/2\right)}\left\{\min_{{\bf c}\in{\cal B}\left(\bar{x}^{\bf A}_{i}/2,\zeta\right)}\left\{e_{i}+c_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf c})\right\}\right\},$ (73)
and
[…]
For any ${\bf c}^{1}\in{\cal B}(\bar{x}^{\bf A}_{i}/2,\zeta)$ and ${\bf c}^{2}\in{\cal B}(\bar{x}^{\bf A}_{i}/2,\bar{\zeta}-\zeta)$, we have ${\bf 1}^{\top}{\bf c}^{1}={\bf 1}^{\top}{\bf c}^{2}$ and $c_{i}^{1}=c_{i}^{2}$. Also, note that
$\mu_{j}\left(\frac{\bar{x}^{\bf A}_{i}}{2},\zeta\right)$ $=\left(\zeta-\bar{s}^{\bf A}_{i,j}-\frac{\bar{x}^{\bf A}_{i}}{2}\right)^{+}\left(Q(\zeta)-Q(\bar{\zeta}-\zeta)\right)+\bar{x}^{\bf A}_{j}Q(\bar{\zeta}-\zeta)$
$=\left(\bar{x}^{\bf A}_{j}-(\bar{\zeta}-\zeta)+\frac{\bar{x}^{\bf A}_{i}}{2}\right)^{+}\left(Q(\zeta)-Q(\bar{\zeta}-\zeta)\right)+\bar{x}^{\bf A}_{j}Q(\bar{\zeta}-\zeta)$
$=\left(\bar{x}^{\bf A}_{j}\wedge\left(\bar{\zeta}-\zeta-\frac{\bar{x}^{\bf A}_{i}}{2}\right)\right)\left(Q(\bar{\zeta}-\zeta)-Q(\zeta)\right)+\bar{x}^{\bf A}_{j}Q(\zeta)$
$=\nu_{j}\left(\frac{\bar{x}^{\bf A}_{i}}{2},\bar{\zeta}-\zeta\right).$
Thus, ${\cal B}(\bar{x}^{\bf A}_{i}/2,\zeta)={\cal B}(\bar{x}^{\bf A}_{i}/2,\bar{\zeta}-\zeta)$ by definition. Hence, without loss of generality, we assume that
$\tilde{\zeta}\leq\bar{\zeta}/2.$ (75)
Fix $x\in[0,\bar{x}^{\bf A}_{i}/2)$ and let
$\zeta^{*}=\begin{cases}\tilde{\zeta},&\text{if }0\leq\tilde{\zeta}\leq(\bar{\zeta}-\bar{x}^{\bf A}_{i})/2+x;\\ (\bar{\zeta}-\bar{x}^{\bf A}_{i})/2+x,&\text{if }(\bar{\zeta}-\bar{x}^{\bf A}_{i})/2+x<\tilde{\zeta}\leq\bar{\zeta}/2.\end{cases}$ (76)
Since $\tilde{\zeta}\in{\cal Z}(\bar{x}^{\bf A}_{i}/2)$ by (73), it is easy to check that
$x+\left(\tilde{\zeta}-\frac{\bar{x}^{\bf A}_{i}}{2}\right)\leq\zeta^{*}\leq\left(\frac{\bar{\zeta}-\bar{x}^{\bf A}_{i}}{2}+x\right)\wedge\tilde{\zeta}$. (77)
We next define
[…]
and
$\eta^{*}\coloneqq\zeta^{*}Q(\zeta^{*})+(\bar{\zeta}-\zeta^{*})Q(\bar{\zeta}-\zeta^{*}).$ (79)
Since $\zeta^{*}\leq\tilde{\zeta}\leq\bar{\zeta}-\tilde{\zeta}\leq\bar{\zeta}-\zeta^{*}$ and the mapping $x\mapsto xQ(x)$ is concave by (21c), we obtain
$\eta^{*}\leq\tilde{\eta}$. (80)
[…]
$\mu_{j}(x,\zeta^{*})\leq\mu_{j}\left(\frac{\bar{x}^{\bf A}_{i}}{2},\tilde{\zeta}\right),\quad\text{for all }j\neq i,l,$ (81)
since $\zeta^{*}\leq\tilde{\zeta}$ and $Q(x)$ is decreasing in $x$ by Assumption 1(ii). Further, we claim that
$\nu_{j}\left(\frac{\bar{x}^{\bf A}_{i}}{2},\tilde{\zeta}\right)\leq\nu_{j}(x,\zeta^{*}),\quad\text{for all }j\neq i.$ (82)
For each $j\neq i$, we observe the following relationship
[…]
Define
$d_{i}\coloneqq xQ(\zeta^{*})+(\bar{x}^{\bf A}_{i}-x)Q(\bar{\zeta}-\zeta^{*}),$ (83)
and
$d_{l}\coloneqq\mu_{l}(x,\zeta^{*})\vee\tilde{c}_{l}.$ (84)
Then by (79) and (83), we can obtain that
$\eta^{*}-d_{i}$ (85)
$=\left(\zeta^{*}-x\right)\left(Q(\zeta^{*})-Q(\bar{\zeta}-\zeta^{*})\right)+(\bar{\zeta}-\bar{x}^{\bf A}_{i})Q(\bar{\zeta}-\zeta^{*})$
$=\left(\left(\zeta^{*}-\bar{s}^{\bf A}_{i,l}-x\right)^{+}+\bar{s}^{\bf A}_{i,l}\wedge\left(\zeta^{*}-x\right)\right)\left(Q(\zeta^{*})-Q(\bar{\zeta}-\zeta^{*})\right)+\left(\bar{x}^{\bf A}_{l}+\bar{s}^{\bf A}_{i,l}\right)Q(\bar{\zeta}-\zeta^{*})$
$=\mu_{l}(x,\zeta^{*})+\bar{s}^{\bf A}_{i,l}Q(\bar{\zeta}-\zeta^{*})+\left(\bar{s}^{\bf A}_{i,l}\wedge\left(\zeta^{*}-x\right)\right)\left(Q(\zeta^{*})-Q(\bar{\zeta}-\zeta^{*})\right)$
$\leq\mu_{l}(x,\zeta^{*})+\bar{s}^{\bf A}_{i,l}Q(\bar{\zeta}-\zeta^{*})+\sum_{j\neq i,l}\left(\bar{x}^{\bf A}_{j}\wedge\left(\zeta^{*}-x\right)\right)\left(Q(\zeta^{*})-Q(\bar{\zeta}-\zeta^{*})\right)$
$=\mu_{l}(x,\zeta^{*})+\sum_{j\neq i,l}\nu_{j}(x,\zeta^{*}).$
Furthermore, since
$\sum_{j\neq i,l}(\zeta^{*}-\bar{s}^{\bf A}_{i,j}-x)^{+}$ $\leq\sum_{j\neq i,l}\left(\hat{\zeta}-\bar{s}^{\bf A}_{i,j}-x\right)^{+}+\zeta^{*}-\hat{\zeta}=\zeta^{*}-x-\bar{x}^{\bf A}_{l}$
and $\hat{\zeta}=\bar{\zeta}-\bar{x}^{\bf A}_{i}+x\geq\zeta^{*}$, we have
$\eta^{*}-d_{i}$ (86)
$=\left(\zeta^{*}-x\right)\left(Q(\zeta^{*})-Q(\bar{\zeta}-\zeta^{*})\right)+(\bar{\zeta}-\bar{x}^{\bf A}_{i})Q(\bar{\zeta}-\zeta^{*})$
$=\left[\bar{x}^{\bf A}_{l}\wedge(\zeta^{*}-x)+\left(\zeta^{*}-x-\bar{x}^{\bf A}_{l}\right)^{+}\right]\left(Q(\zeta^{*})-Q(\bar{\zeta}-\zeta^{*})\right)+\left(\bar{x}^{\bf A}_{l}+\bar{s}^{\bf A}_{i,l}\right)Q(\bar{\zeta}-\zeta^{*})$
$=\nu_{l}(x,\zeta^{*})+\bar{s}^{\bf A}_{i,l}Q(\bar{\zeta}-\zeta^{*})+\left(\zeta^{*}-x-\bar{x}^{\bf A}_{l}\right)^{+}\left(Q(\zeta^{*})-Q(\bar{\zeta}-\zeta^{*})\right)$
$\geq\nu_{l}(x,\zeta^{*})+\bar{s}^{\bf A}_{i,l}Q(\bar{\zeta}-\zeta^{*})+\sum_{j\neq i,l}\left(\zeta^{*}-\bar{s}^{\bf A}_{i,j}-x\right)^{+}\left(Q(\zeta^{*})-Q(\bar{\zeta}-\zeta^{*})\right)$
$=\nu_{l}(x,\zeta^{*})+\sum_{j\neq i,l}\mu_{j}(x,\zeta^{*}).$
Step 4. Recall that $d_{i}$ and $d_{l}$ are defined in (83) and (84), respectively. We define
$\Delta\coloneqq\eta^{*}-d_{i}-d_{l}-\sum_{j\neq i,l}\tilde{c}_{j},$ (87)
and for all $j\neq i,l$,
$d_{j}\coloneqq\begin{cases}\tilde{c}_{j}+\dfrac{\nu_{j}(x,\zeta^{*})-\tilde{c}_{j}}{\sum_{k\neq i,l}(\nu_{k}(x,\zeta^{*})-\tilde{c}_{k})}\Delta,&\text{if }\Delta>0;\\[15.0pt] \tilde{c}_{j}+\dfrac{\tilde{c}_{j}-\mu_{j}(x,\zeta^{*})}{\sum_{k\neq i,l}(\tilde{c}_{k}-\mu_{k}(x,\zeta^{*}))}\Delta,&\text{otherwise}.\end{cases}$ (88)
We first claim that
${\bf d}=(d_{1},\dots,d_{m})^{\top}\in{\cal B}(x,\zeta^{*}).$ (89)
Observe that
${\bf 1}^{\top}{\bf d}=\sum_{j\neq i,l}d_{j}+d_{i}+d_{l}=\sum_{j\neq i,l}\tilde{c}_{j}+\Delta+d_{i}+d_{l}=\eta^{*}.$
[…]
$\tilde{c}_{l}\leq\nu_{l}(x,\zeta^{*}),$ (90)
and
$\mu_{j}(x,\zeta^{*})\leq\tilde{c}_{j}\leq\nu_{j}(x,\zeta^{*}),\quad\text{for all }j\neq i,l.$ (91)
Additionally, by (84), (85), (86), and (90), it is straightforward to see that
$\begin{cases}\Delta\leq\sum_{k\neq i,l}(\nu_{k}(x,\zeta^{*})-\tilde{c}_{k}),&\text{if }\Delta>0;\\ \Delta\geq-\sum_{k\neq i,l}(\tilde{c}_{k}-\mu_{k}(x,\zeta^{*})),&\text{otherwise.}\end{cases}$ (92)
Accordingly, $\mu_{j}(x,\zeta^{*})\leq d_{j}\leq\nu_{j}(x,\zeta^{*})$ for all $j\neq i$, and hence, (89) holds.
Next, we want to show that
$\tilde{c}_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}(\tilde{{\bf c}})>d_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf d}).$ (93)
To that end, we use the following lemma whose proof is deferred to Appendix C.
Lemma 4.
For fixed $i\in[m]$, let $l=\mathop{\arg\max}_{j\neq i}\bar{x}^{\bf A}_{j}$. Define $\tilde{{\bf c}}$, $d_{i}$, and $d_{l}$ as in (74), (83), and (84), respectively. If (76) holds, then $\tilde{c}_{i}-d_{i}>0$ and $\tilde{c}_{i}-d_{i}\geq d_{l}-\tilde{c}_{l}.$
If $\boldsymbol{\pi}(\tilde{{\bf c}})=\boldsymbol{\pi}({\bf d})$, (93) holds by Lemma 4. Otherwise, let $\hat{{\bf c}}\coloneqq{\bf d}\wedge\tilde{{\bf c}}$. Then, $\hat{c}_{l}=\tilde{c}_{l}$ by (84), $\hat{c}_{i}=d_{i}$ by Lemma 4, and for all $j\neq i,l$,
$\hat{c}_{j}=\begin{cases}\tilde{c}_{j},\quad\quad&\text{if }\Delta>0;\\ d_{j},\quad\quad&\text{otherwise.}\end{cases}$ (94)
by (88) and (91). Therefore, we have
$\tilde{c}_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}(\tilde{{\bf c}})-\left(d_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf d})\right)$
$=\tilde{c}_{i}-d_{i}+{\bf A}_{i}^{\top}\left(\boldsymbol{\pi}(\tilde{{\bf c}})-\boldsymbol{\pi}({\bf d})\right)$
$\geq\tilde{c}_{i}-d_{i}+{\bf A}_{i}^{\top}\left(\boldsymbol{\pi}(\hat{{\bf c}})-\boldsymbol{\pi}({\bf d})\right)$ (95)
$>\tilde{c}_{i}-d_{i}-\left\|\boldsymbol{\pi}(\hat{{\bf c}})-\boldsymbol{\pi}({\bf d})\right\|_{1}$ (96)
$\geq\tilde{c}_{i}-d_{i}-\left\|\hat{{\bf c}}-{\bf d}\right\|_{1}$ (97)
$=\tilde{c}_{i}-d_{i}+\sum_{j\neq i}\left(\hat{c}_{j}-d_{j}\right),$
$=\begin{cases}\tilde{\eta}-\eta^{*},\quad\quad&\text{if }\Delta>0;\\ \tilde{c}_{i}-d_{i}-(d_{l}-\tilde{c}_{l}),\quad\quad&\text{otherwise},\end{cases}$ (98)
where (95) holds since ${\bf A}_{i}$ is nonnegative and $\boldsymbol{\pi}(\cdot)$ is increasing (see Eisenberg and Noe, 2001, Lemma 5), (96) is satisfied since $A_{ki}\leq\sum_{j=1}^{m}A_{kj}<1$ for all $k\in[m]$, (97) stems from the fact that $\boldsymbol{\pi}(\cdot)$ is $\ell_{1}$-nonexpansive (see Eisenberg and Noe, 2001, Lemma 5), and (98) follows since we have ${\bf 1}^{\top}\tilde{{\bf c}}=\tilde{\eta}$ and ${\bf 1}^{\top}{\bf d}=\eta^{*}$. Consequently, by (80) and Lemma 4, (93) follows.
Note that $\tilde{\zeta}\geq\bar{x}^{\bf A}_{i}/2$ since $\tilde{\zeta}\in{\cal Z}(\bar{x}^{\bf A}_{i}/2)$. Thus, $\zeta^{*}\in{\cal Z}(x)$ by (77), and we finally have
[…]
$>e_{i}+d_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf d})$
$\geq\min_{\zeta\in{\cal Z}(x)}\min_{{\bf c}\in{\cal B}(x,\zeta)}\left\{e_{i}+c_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf c})\right\}$ (100)
$=\tilde{g}_{i}(x),$ (101)
where (99) stems from (71), (73) and (74), and (100) holds since $\zeta^{*}\in{\cal Z}(x)$ and ${\bf d}\in{\cal B}(x,\zeta^{*})$. Then, by (71), we have (66), and the proof is complete. ∎
Appendix C Proofs for Auxiliary Results
Proof.
Proof of Lemma 3. Assume by contradiction that $\bar{p}^{\bf A}>p_{2}^{\bf A}$. For $k=0,1,2,\ldots,$ define
$\rho_{k+1}^{\dagger}=Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-{\bf A}^{\top}\boldsymbol{\lambda}_{k}^{\dagger}\right)^{+}}{\rho_{k}^{\dagger}}\wedge z_{j}\right)~~\text{and}~~\boldsymbol{\lambda}_{k+1}^{\dagger}={\bf L}\wedge\left({\bf e}+{\bf z}\rho_{k}^{\dagger}+{\bf A}^{\top}\boldsymbol{\lambda}_{k}^{\dagger}\right),$ (102)
where
$\rho_{0}^{\dagger}\coloneqq Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge z_{j}\right)~~\text{and}~~\boldsymbol{\lambda}_{0}^{\dagger}\coloneqq{\bf L}\wedge\left({\bf e}+{\bf z}\rho_{0}^{\dagger}+{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right).$ (103)
Fix ${\bf x}\leq\bar{{\bf x}}^{\bf A}$. Then, $p_{1}\coloneqq Q({\bf 1}^{\top}{\bf x})\geq Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})=\bar{p}^{\bf A}>p_{2}^{\bf A}$. Moreover, by (15) and Assumption 1(ii), we have
$p_{2}^{\bf A}$ $=Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-x_{j}p_{1}-{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)\right)$
$\geq Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge z_{j}\right)$
$=\rho_{0}^{\dagger}.$
Consequently, we observe that $\boldsymbol{\pi}^{\bf A}={\bf L}\wedge\left({\bf e}+{\bf x}p_{1}+({\bf z}-{\bf x})p_{2}^{\bf A}+{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right)\geq{\bf L}\wedge\left({\bf e}+{\bf z}\rho_{0}^{\dagger}+{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right)=\boldsymbol{\lambda}_{0}^{\dagger}$,
If $\rho_{k}^{\dagger}\leq p_{2}^{\bf A}$ and $\boldsymbol{\lambda}_{k}^{\dagger}\leq\boldsymbol{\pi}^{\bf A}$ for some $k\geq 0$, then
$\rho_{k+1}^{\dagger}=Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-{\bf A}^{\top}\boldsymbol{\lambda}_{k}^{\dagger}\right)^{+}}{\rho_{k}^{\dagger}}\wedge z_{j}\right)\leq Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-x_{j}Q({\bf 1}^{\top}{\bf x})-{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)\right)=p_{2}^{\bf A}$ (104)
and thus
$\boldsymbol{\lambda}_{k+1}^{\dagger}={\bf L}\wedge\left({\bf e}+{\bf z}\rho_{k}^{\dagger}+{\bf A}^{\top}\boldsymbol{\lambda}_{k}^{\dagger}\right)\leq{\bf L}\wedge\left({\bf e}+{\bf x}p_{1}+({\bf z}-{\bf x})p_{2}^{\bf A}+{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right)=\boldsymbol{\pi}^{\bf A}.$ (105)
Hence, by induction, we obtain $\rho_{k}^{\dagger}\leq p_{2}^{\bf A}$ and $\boldsymbol{\lambda}_{k}\leq\boldsymbol{\pi}^{\bf A}$ for all $k$, suggesting that $\bar{p}^{\bf A}\leq p_{2}^{\bf A}$ and $\bar{\boldsymbol{\pi}}^{\bf A}\leq\boldsymbol{\pi}^{\bf A}$ since $\rho_{k}^{\dagger}\rightarrow\bar{p}^{\bf A}$ and $\boldsymbol{\lambda}_{k}^{\dagger}\rightarrow\bar{\boldsymbol{\pi}}^{\bf A}$ as $k$ increases. This contradicts the assumption and indicates that $\bar{p}^{\bf A}\leq p_{2}^{\bf A}$. ∎
Proof.
Proof of Lemma 4. We first let $\delta=\tilde{\zeta}-\zeta^{*}$. Then, by (76),
$0\leq\delta\leq\frac{\bar{x}^{\bf A}_{i}}{2}-x,$ (106)
and we observe that
$\tilde{c}_{i}-d_{i}$ $=\frac{\bar{x}^{\bf A}_{i}}{2}\left(Q(\tilde{\zeta})+Q(\bar{\zeta}-\tilde{\zeta})\right)-xQ(\zeta^{*})-(\bar{x}^{\bf A}_{i}-x)Q(\bar{\zeta}-\zeta^{*})$ (107)
$=\left(x+\delta\right)Q\left(\zeta^{*}+\delta\right)-xQ\left(\zeta^{*}\right)-\left(\bar{x}^{\bf A}_{i}-x\right)Q\left(\bar{\zeta}-\zeta^{*}\right)$
[…]
$>\delta\bigg(f_{x}(x+\theta_{1},\zeta^{*}-x)\bigg.-f_{x}(\bar{x}^{\bf A}_{i}-x-\theta_{1},\bar{\zeta}-\zeta^{*}-\bar{x}^{\bf A}_{i}+x)\bigg)$
$\geq 0,$
where the first inequality stems from (21c) and (106), and the second inequality is satisfied since $\zeta^{*}-x\leq\bar{\zeta}-\zeta^{*}-\bar{x}^{\bf A}_{i}+x$ and (21d) holds. This proves the first statement of the lemma.
[…]
$=\tilde{\vartheta}\left(Q\left(\tilde{\zeta}\right)-Q\left(\bar{\zeta}-\tilde{\zeta}\right)\right)+\left(\bar{x}^{\bf A}_{i}+\left(\bar{x}^{\bf A}_{l}-\bar{s}^{\bf A}_{i,l}\right)^{+}\right)\left(Q\left(\bar{\zeta}-\tilde{\zeta}\right)-Q\left(\bar{\zeta}-\zeta^{*}\right)\right)$
$\qquad-\vartheta^{*}\left(Q\left(\zeta^{*}\right)-Q\left(\bar{\zeta}-\zeta^{*}\right)\right)+\left(\bar{x}^{\bf A}_{l}\wedge\bar{s}^{\bf A}_{i,l}\right)\left(Q\left(\bar{\zeta}-\tilde{\zeta}\right)-Q\left(\bar{\zeta}-\zeta^{*}\right)\right)$
$=\left(\vartheta^{*}+\delta\right)\left(Q\left(\zeta^{*}+\delta\right)-Q\left(\bar{\zeta}-\zeta^{*}-\delta\right)\right)+\left(\bar{x}^{\bf A}_{i}+\left(\bar{x}^{\bf A}_{l}-\bar{s}^{\bf A}_{i,l}\right)^{+}\right)\left(Q\left(\bar{\zeta}-\tilde{\zeta}\right)-Q\left(\bar{\zeta}-\zeta^{*}\right)\right)$
[…]
$\qquad+\left(\bar{x}^{\bf A}_{l}\wedge\bar{s}^{\bf A}_{i,l}\right)\left(Q\left(\bar{\zeta}-\tilde{\zeta}\right)-Q\left(\bar{\zeta}-\zeta^{*}\right)\right)$
$\geq\delta\Big\{Q\left(\zeta^{*}+\theta_{2}\right)-Q\left(\bar{\zeta}-\zeta^{*}-\theta_{2}\right)-\left(\bar{x}^{\bf A}_{i}+\left(\bar{x}^{\bf A}_{l}-\bar{s}^{\bf A}_{i,l}\right)^{+}\right)Q^{\prime}\left(\bar{\zeta}-\zeta^{*}-\theta_{2}\right)$
$\qquad+\left(\vartheta^{*}+\theta_{2}\right)\left(Q^{\prime}\left(\zeta^{*}+\theta_{2}\right)+Q^{\prime}\left(\bar{\zeta}-\zeta^{*}-\theta_{2}\right)\right)\Big\}$
$=\delta\left\{f_{x}(\vartheta^{*}+\theta_{2},\zeta^{*}-\vartheta^{*})-f_{x}\left(\bar{x}^{\bf A}_{i}+\left(\bar{x}^{\bf A}_{l}-\bar{s}^{\bf A}_{i,l}\right)^{+}-\vartheta^{*}-\theta_{2},\bar{\zeta}-\zeta^{*}-\bar{x}^{\bf A}_{i}-\left(\bar{x}^{\bf A}_{l}-\bar{s}^{\bf A}_{i,l}\right)^{+}+\vartheta^{*}\right)\right\}$
$>\delta\Big\{f_{x}\left(\bar{x}^{\bf A}_{i}+\left(\bar{x}^{\bf A}_{l}-\bar{s}^{\bf A}_{i,l}\right)^{+}-\vartheta^{*}-\theta_{2},\zeta^{*}-\vartheta^{*}\right)$
$\qquad-f_{x}\left(\bar{x}^{\bf A}_{i}+\left(\bar{x}^{\bf A}_{l}-\bar{s}^{\bf A}_{i,l}\right)^{+}-\vartheta^{*}-\theta_{2},\bar{\zeta}-\zeta^{*}-\bar{x}^{\bf A}_{i}-\left(\bar{x}^{\bf A}_{l}-\bar{s}^{\bf A}_{i,l}\right)^{+}+\vartheta^{*}\right)\Big\}$
$=0,$
[…]
$2\left((\zeta^{*}-\bar{s}^{\bf A}_{i,l})\vee x\right)\leq 2\left\{\left(\frac{\bar{\zeta}-\bar{x}^{\bf A}_{i}}{2}+x-\bar{s}^{\bf A}_{i,l}\right)\vee x\right\}\leq\left(\bar{x}^{\bf A}_{l}-\bar{s}^{\bf A}_{i,l}\right)^{+}+2x<\left(\bar{x}^{\bf A}_{l}-\bar{s}^{\bf A}_{i,l}\right)^{+}+\bar{x}^{\bf A}_{i}-2\theta_{2},$
and thus, the second inequality of (108) follows from (21c). Similarly, we obtain
$2\left(\zeta^{*}-\left(\zeta^{*}-\bar{s}^{\bf A}_{i,l}\right)\vee x\right)=\left(2\bar{s}^{\bf A}_{i,l}\right)\wedge(2\zeta^{*}-2x)=\left(2\bar{s}^{\bf A}_{i,l}\right)\wedge(\bar{\zeta}-\bar{x}^{\bf A}_{i})=\bar{\zeta}-\bar{x}^{\bf A}_{i}-(\bar{x}^{\bf A}_{l}-\bar{s}^{\bf A}_{i,l})^{+},$
which implies that the last equality of (108) holds. This completes the proof. ∎